### v1

# Mingle: Mixtures of Null-Space Gated Low-Rank Experts for Test-Time Continual Model Merging

Abstract
Continual model merging integrates independently fine-tuned models sequentially without access to original training data, providing a scalable and efficient solution to continual learning. However, current methods still face critical challenges, notably parameter interference among tasks and limited adaptability to evolving test distributions. The former causes catastrophic forgetting of integrated tasks, while the latter hinders effective adaptation to new tasks. To address these, we propose Mingle, a novel framework for test-time continual model merging, which leverages test-time adaptation using a small set of unlabeled test samples from the current task to dynamically guide the merging process. Mingle employs a mixture-of-experts architecture composed of parameter-efficient, low-rank experts, enabling efficient adaptation and improving robustness to distribution shifts. To mitigate catastrophic forgetting, we propose Null-Space Constrained Gating, which restricts gating updates to subspaces orthogonal to prior task representations. This suppresses activations on old task inputs and preserves model behavior on past tasks. To further balance stability and adaptability, we design an Adaptive Relaxation Strategy, which dynamically adjusts the constraint strength based on interference signals captured during test-time adaptation. Extensive experiments on standard continual merging benchmarks demonstrate that Mingle achieves robust generalization, reduces forgetting significantly, and consistently surpasses previous state-of-the-art methods by 7-9% on average across diverse task orders.
1 Introduction
Continual learning seeks to incrementally adapt machine learning models to new tasks without forgetting previously learned knowledge, addressing the critical challenge of catastrophic forgetting [41]. However, conventional continual learning approaches typically require continuous access to original training data, raising significant concerns about privacy and substantial computational overhead due to retraining efforts, thus limiting their applicability in dynamic, data-sensitive environments.
To address these limitations, recent works have explored an alternative paradigm known as continual model merging (CMM), which sequentially integrates independently fine-tuned models directly in parameter space, without revisiting any past training data [37, 49, 68]. CMM typically operates under a "merge-to-transfer" paradigm: given a pretrained model $\theta_{0}$ and independently fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$, a unified model is constructed by combining task-specific weight updates $\Delta\theta_{t}=\theta_{t}-\theta_{0}$ via weighted averaging or projection-based strategies [23, 77, 67].
Despite its advantages in scalability, data privacy, and distributed training capabilities [42, 13, 55], existing CMM methods still encounter critical issues, notably severe parameter interference between tasks and limited adaptability to evolving test distributions. This parameter interference arises because, as fine-tuned models are incrementally merged, overlapping or conflicting parameter updates accumulate, resulting in severe forgetting of previously learned tasks. To mitigate this interference, recent methods introduce structural constraints such as orthogonal projection [68, 74], model linearization [37, 67], and pruning-based sparsification [82, 86]. However, their effectiveness diminishes as task count grows and interference becomes increasingly entangled. Moreover, models merged across tasks often fail to generalize effectively, particularly when facing unseen or shifting task conditions. These flaws result in severe forgetting of earlier tasks and substantial performance gaps compared to the upper bound achieved by individually fine-tuned models. As shown in Fig. 1, TA [23] suffers from large performance gaps and strong forgetting, reflected by low accuracy and negative backward transfer (BWT). OPCM [68] improves over TA via orthogonalized merging but still shows notable degradation.
Figure 1: Performance after eight task continual merging: accuracy on first four tasks and overall BWT.
To overcome these limitations, we propose a novel continual merging paradigm—Test-Time Continual Model Merging (TTCMM)—which explicitly introduces the concept of test-time adaptation (TTA) [72, 61] into model merging. Unlike prior TTA-based multi-task merging methods [83, 66, 84], which assume simultaneous availability of models and test data from all tasks, TTCMM utilizes only a small set of unlabeled samples from the current task, making it uniquely suited for realistic continual scenarios where revisiting historical data is often infeasible.
In this paper, we propose Mingle (MIxtures of Null-Space Gated Low-rank Experts), a method designed to continually merge independently fine-tuned models at test-time while preserving prior knowledge. Mingle employs a mixture-of-experts architecture [26, 43] composed of lightweight LoRA-based [20] experts, enabling efficient and flexible test-time adaptation. To robustly prevent interference from previously learned tasks, we introduce a novel Null-Space Constrained Gating mechanism, restricting gating updates to task-orthogonal subspaces. Additionally, we propose an Adaptive Relaxation Strategy to dynamically modulate constraint strength based on test-time interference feedback during adaptation.
Extensive experiments on standard continual learning benchmarks show that Mingle consistently outperforms previous state-of-the-art approaches by 7–9% on average, achieving robust generalization and strong resistance to catastrophic forgetting across diverse continual learning scenarios. Remarkably, these improvements are achieved entirely without any access to original training data, demonstrating the effectiveness of our TTCMM paradigm and the power of test-time adaptation in continual learning.
Our contributions are summarized as follows:
  • We formalize test-time continual model merging (TTCMM), a novel framework leveraging unlabeled test samples for merging independently fine-tuned models.
  • We propose Mingle, a TTCMM method with Adaptive Null-Space Constrained Gating to effectively balance stability and plasticity.
  • Extensive experiments show that Mingle achieves state-of-the-art performance, consistently outperforming prior methods in accuracy, robustness and resistance to forgetting.
2 Related Work
Continual Learning.  Continual learning (CL) aims to mitigate catastrophic forgetting [41], where models forget prior tasks when learning new ones. Regularization-based methods constrain updates via importance weights [30, 88, 2, 29, 78], while distillation techniques preserve knowledge by aligning outputs or intermediate features [19, 11, 58, 50]. Replay-based methods store exemplars or generate surrogates using prompts, prototypes, or generators [53, 38, 75, 59, 51]. Dynamic architectures expand capacity through growth or ensembling [34, 94, 40]. Recent work integrates lightweight adapters or prompts into pre-trained models to enhance transfer with minimal updates [85, 22, 76]. CL has also been approached through model merging [77, 23, 27, 66]. One line remains close to conventional CL, incrementally fine-tuning models and merging them to reduce forgetting [39, 40, 15]. These methods typically require access to training data and involve sequential task-wise fine-tuning. In contrast, the continual model merging paradigm [27, 37, 5, 49, 68] merges independently fine-tuned models without revisiting prior data, offering greater scalability and privacy.
Model Merging.  Early work merged models via direct parameter averaging [70, 56], later refined by linear mode connectivity [12, 1]. Task Arithmetic (TA) [23] generalizes this by viewing models as task vectors to be summed. However, TA relies on weight disentanglement [47], which standard fine-tuning often violates, prompting structured training [27, 63]. Beyond averaging, interference-aware methods reweight or sparsify parameters [82, 86], or fuse models via distillation and clustering [83, 73]. LoRA-based tuning [20] introduces additional challenges due to parameter entanglement, spurring gradient-free or retrieval-based fusion strategies [21, 90, 91].
Figure 2: Comparison of three continual learning paradigms. (a) Conventional Continual Learning trains models sequentially with data arriving in stages, without access to previous task data. (b) Continual Model Merging continually fuses independently trained models, without access to any training data. (c) Test-Time Continual Model Merging improves merging by leveraging a few unlabeled test samples from the current task.
Test-Time Adaptation.  TTA updates models at inference to counter test-time distribution shift. Early methods employed self-supervised objectives [64], entropy minimization [72], and regularized updates [61]. Online TTA adapts continuously [24], while batch-wise variants ignore temporal structure [16]. To improve stability, recent methods incorporate confidence filtering [46], EMA [10], partial updates [87], test-time augmentation [89], and adaptive BatchNorm [54]. In vision–language models, TTA leverages prompts or adapters [57, 14, 35]. Our approach draws on TTA to steer merging dynamics, aligning fused models with continuously evolving test distributions.
Discussion.   Most related to our method are MoE-Adapter [85] and WEMOE [66], both leveraging MoE architectures [26, 28]. MoE-Adapter follows conventional CL, where expert modules are embedded in the model and trained as a whole across tasks. In contrast, we adopt a model merging paradigm: experts are extracted from independently fine-tuned models and inserted without training. While WEMOE incorporates TTA, it is designed for multitask learning rather than continual scenarios, relying on simultaneous access to all models and data. By contrast, Mingle is explicitly tailored for the more challenging continual merging setup.
3 Mingle: Mixtures of Null-Space Gated Low-Rank Experts
3.1 Preliminaries
Problem Setting.
We study continual learning in a model merging setting, where a sequence of task-specific models $\{\theta_{1},\dots,\theta_{T}\}$ are independently fine-tuned from a shared pre-trained model $\theta_{0}$, each using a dataset $\mathcal{D}_{i}=\{(x_{j}^{(i)},y_{j}^{(i)})\}$ with label space $\mathcal{C}_{i}\subset\mathcal{Y}$. The goal is to construct a unified model $\theta^{\text{merged}}_{T}$ that generalizes across the combined label space $\mathcal{C}_{1:T}=\bigcup_{i=1}^{T}\mathcal{C}_{i}$.
Unlike conventional continual learning, we assume no access to training data during merging. All adaptation happens directly in parameter space. This paradigm is relevant in scenarios where only final fine-tuned models are retained, while original training data is discarded due to privacy, storage, or accessibility constraints.
To contextualize this, we compare with two related paradigms in Fig. 2:
  • Conventional Continual Learning. A single model $\theta$ is sequentially updated on $\mathcal{D}_{1},\dots,\mathcal{D}_{T}$, discarding previous data. It requires direct training data access and extensive retraining.
  • Continual Model Merging. A sequence of models $\{\theta_{i}\}$ are merged incrementally in parameter space without access to training data and earlier models: $\theta^{\text{merged}}_{t}=\text{Merge}(\theta^{\text{merged}}_{t-1},\theta_{t})$.
  • Test-Time Continual Model Merging. An extension of the above where a small unlabeled subset $\mathcal{D}^{\text{seed}}_{t}\subset\mathcal{D}^{\text{test}}_{t}$ (e.g., 5 samples per class) is available at each stage to provide lightweight task-specific guidance. We refer to $\mathcal{D}^{\text{seed}}_{t}$ as the seed samples of task $t$.
Existing Continual Merging Strategies.
Let $\theta_{0}$ denote the parameters of a pre-trained model. The corresponding task vector is defined as $\Delta\theta_{t}=\theta_{t}-\theta_{0}$.
  • Continual Task Arithmetic (C. TA). A simple additive merge [23]: $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda\Delta\theta_{t}$, where $\lambda$ is a scalar. While training-free, it is sensitive to $\lambda$ and prone to task interference.
  • Orthogonal Projection-based Continual Merging (OPCM). Tang et al. [68] propose projecting each $\Delta\theta_{t}$ onto the orthogonal complement of previous directions: $\theta^{\text{merged}}_{t}=\theta_{0}+\frac{1}{\lambda_{t}}\left[\lambda_{t-1}\Delta\theta^{\text{merged}}_{t-1}+\mathcal{P}^{(t-1)}(\Delta\theta_{t})\right]$, where $\mathcal{P}^{(t-1)}$ retains components orthogonal to previous updates. This reduces interference but ignores adaptation to task distributions.
To address these issues, we present Mingle, which leverages $\mathcal{D}^{\text{seed}}_{t}$ to modulate the integration of $\theta_{t}$ at test-time, enhancing alignment to test distribution and mitigating task interference.
3.2 Motivation and Theoretical Analysis
Most existing continual model merging methods combine fine-tuned models via static averaging, where each expert is assigned fixed coefficients, thereby enforcing the same mixing rule across the whole input space. Consequently, it cannot specialize to regions where one expert is clearly superior. In contrast, a Mixture-of-Experts (MoE) equips every input with a data-dependent gate $g(x)=(g_{1}(x),\dots,g_{T}(x))$ that selects or re-weights experts on-the-fly. We give a formal comparison between static averaging and dynamic MoE under a noisy-routing scenario.  [^1] Symbols and proofs are deferred to App. A
Theorem 1 (Dynamic MoE versus Static Averaging).
Let $\{(D_{t},f_{t})\}_{t=1}^{T}$ be $T$ independent tasks with priors $P(t)$ and per-task risks $R_{t}(i)$. For any static mixture $h_{\mathrm{static}}(x)=\sum_{i=1}^{T}\alpha_{i}\,f_{i}(x)$ and any hard-routed MoE $h_{\mathrm{MoE}}(x)=f_{i^{\star}(x)}(x)$ with task-specific routing errors $\varepsilon_{t}$:
$R(h_{\mathrm{MoE}})\;=\;R_{\mathrm{ideal}}+\sum_{t=1}^{T}P(t)\,\varepsilon_{t}\bigl{(}R_{\text{wrong},t}-R_{t}(t)\bigr{)},$ (1)
where $R_{\mathrm{ideal}}=\sum_{t}P(t)R_{t}(t)$ and $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$. Moreover,
  1. 1.
    (Perfect routing) If $\varepsilon_{t}=0$ for all $t$, then $\inf_{g}R(h_{\mathrm{MoE}})\;<\;\inf_{\boldsymbol{\alpha}}R(h_{\mathrm{static}})$ whenever at least two tasks disagree on their best expert.
  2. 2.
    (Noisy routing)  If $\sum_{t}P(t)\varepsilon_{t}\bigl{(}R_{\text{wrong},t}-R_{t}(t)\bigr{)}<R^{*}_{\mathrm{static}}-R_{\mathrm{ideal}},$ where $R^{*}_{\text{static}}=\inf_{\boldsymbol{\alpha}}R(h_{\text{static}})$, then the MoE still attains lower risk than any static mixture.
The theory above motivates a design that (i) keeps experts specialized and (ii) prevents interference between tasks. Our Mingle framework achieves both goals by combining
  • Low-rank experts $f_{t}$ that capture task-specific variations with minimal parameters, and
  • Null-space constrained gating that projects gradient updates away from subspaces spanned by previously activated features, keeping $\varepsilon_{t}$ small without harming earlier experts.
3.3 Low-Rank Expert Mixtures for Continual Model Merging
We adopt MoE framework for continual model merging, in which each task $i$ is equipped with a low-rank expert $f_{i}$ and an associated input-dependent gating function $g_{i}$. These components are injected into the linear layers of the backbone (e.g., CLIP visual encoder). The gate $g_{i}$ modulates expert activation based on the input features, allowing for fine-grained, localized task specialization.
Mixtures of Low-Rank Expert.  When a new task $t$ arrives, a dedicated expert $f_{t}$ and its gate $g_{t}$ are appended to the model. The output of a given $l$-th layer [^2] The layer index $l$ is omitted hereafter whenever it does not cause ambiguity. can be formulated as follows:
$\theta_{t}^{\text{merged},(l)}(X)=\theta_{t-1}^{\text{merged},(l)}(X)+g_{t}^{(l)}(X)\cdot f_{t}^{(l)}(X)=\theta_{0}^{(l)}(X)+\sum_{i=1}^{t}g_{i}^{(l)}(X)\cdot f_{i}^{(l)}(X).$ (2)
where only the gate $g_{t}$ is adaptable during testing, while all experts $\{f_{i}\}_{i=1}^{t}$ and old gates $\{g_{i}\}_{i=1}^{t-1}$ remain frozen to preserve prior knowledge. To construct expert $f_{t}$, we first project the task vector $\Delta\theta_{t}$ onto the orthogonal complement of previously learned directions, following OPCM [68]:
$\mathcal{P}^{(t-1)}(\Delta\theta_{t})=\sum_{\begin{subarray}{c}p,q=a,p\neq q\end{subarray}}^{m,n}\langle\Delta\theta_{t},u_{p}^{(t-1)}v_{q}^{(t-1)\top}\rangle_{F}u_{p}^{(t-1)}v_{q}^{(t-1)\top},$ (3)
where $u_{p}^{(t-1)}$ and $v_{q}^{(t-1)}$ are the $p$-th and $q$-th singular vectors from the singular value decomposition (SVD) of previous experts $\sum_{i=1}^{t-1}f_{i}(X)$. This projection removes previously learned directions to mitigate interference. We then apply a rank-$r$ truncated SVD for $\mathcal{P}^{(t-1)}(\Delta\theta_{t})$ to construct a low-rank expert [20]:
$f_{t}=BA=(\tilde{U}\tilde{\Sigma})\tilde{V}^{\top},$ (4)
where $\tilde{U}\in\mathbb{R}^{d_{1}\times r}$, $\tilde{\Sigma}\in\mathbb{R}^{r\times r}$, and $\tilde{V}\in\mathbb{R}^{d_{2}\times r}$, retaining the top $r$ singular components. The resulting expert captures the principal directions while significantly reducing parameter overhead.
Each gating function is implemented as a linear projection:
$g_{t}(X)=W_{t}^{(g)\top}X+b_{t}^{(g)},$ (5)
where $W_{t}^{(g)}\in\mathbb{R}^{d\times 1}$ and $b_{t}^{(g)}\in\mathbb{R}$ are learnable parameters. The gating function is adapted at test time using a small number of unlabeled test data.
Test-Time Adaptation.  Given class text embeddings $T=[T_{1},\dots,T_{|\mathcal{C}_{t}|}]\in\mathbb{R}^{|\mathcal{C}_{t}|\times d}$, where each $T_{i}$ is a prompt (e.g., “a photo of a dog”), the model computes logits via scaled cosine similarity: $p(y\mid x)=\mathrm{softmax}(z(x)),\ z(x)=\tau\cdot\mathrm{norm}(\phi_{\text{img}}(x))\cdot\mathrm{norm}(\phi_{\text{text}}(T))^{\top},$ where $\phi_{\text{img}}$ and $\phi_{\text{text}}$ are the image and text encoders, $\mathrm{norm}(\cdot)$ denotes $L_{2}$-normalization, and $\tau$ is a temperature scalar.
To encourage the merged model to retain task-specific behavior, we minimize the Kullback–Leibler divergence between its prediction and that of the corresponding individual fine-tuned model $\theta_{t}$. We define the adaptation objective as:
$L_{t}=\mathbb{E}_{x\sim\mathcal{D}^{\text{seed}}_{t}}\left[\mathrm{KL}\left(p(x;\;\theta^{\text{merged}}_{t})\,\|\,p(x;\;\theta_{t})\right)\right].$ (6)
3.4 Adaptive Null-Space Constrained Gating for Interference Mitigation
When merging models continually, the primary challenge of gating is to integrate new experts without disturbing prior task predictions. Consider two experts $f_{1},f_{2}$ and their corresponding gates $g_{1},g_{2}$. When evaluating on the first task domain $X_{1}$, the interference from $g_{2}$ can be quantified as:
$\xi(g_{2})=\left\|g_{1}(X_{1})\cdot f_{1}(X_{1})+g_{2}(X_{1})\cdot f_{2}(X_{1})-f_{1}(X_{1})\right\|^{2}.$ (7)
This measures the deviation introduced by $g_{2}$ on the domain where $f_{1}$ originally dominates. A desirable gating function should suppress $g_{2}(X_{1})$, ensuring predictions on $X_{1}$ remain unaffected. However, as $X_{1}$ becomes inaccessible after adaptation, this error becomes unobservable and cannot be minimized directly, resulting in prediction drift and catastrophic forgetting.
Hard Null-Space Projection.  After finishing task $t$ we cache $l$-th layer activations on a seed buffer $\mathcal{D}_{t}^{\text{seed}}$ and estimate its covariance $C_{t}^{(l)}=\mathbb{E}[\theta_{t}^{\mathrm{merged},(l)}(X_{t})\,\theta_{t}^{\mathrm{merged},(l)}(X_{t})^{\top}]\in\mathbb{R}^{d\times d}$. Truncated SVD yields the top-$k$ dominant subspaces $\tilde{U}_{t}^{(l)}\in\mathbb{R}^{d\times k}$. Concatenating all previous tasks and orthonormalising: $U_{t}^{(l)}=\operatorname{orthonorm}\bigl{[}U_{t-1}^{(l)}|\tilde{U}_{t}^{(l)}\bigr{]}\in\mathbb{R}^{d\times tk}$, the hard projector is $P_{t}=I-U_{t}^{(l)}\,U_{t}^{(l)\top}\in\mathbb{R}^{d\times d}.$ To eliminate interference on task $\leq t-1$, the gate update for task $t$ is $W_{t}^{(g,l)}\leftarrow W_{t}^{(g,l)}-\eta\,P_{t-1}^{(l-1)}\nabla L_{t}^{(l)},$ which may also removes gradient components that are informative for task $t$ whenever its feature support overlaps $\operatorname{span}(U_{t-1})$.
Input: pre-trained model $\theta_{0}$ and fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$; seed data $\{\mathcal{D}^{\text{seed}}_{t}\}_{t=1}^{T}$; hyper-parameters $k,\beta,\gamma$; learning rate $\eta$
Output: Merged model $\theta_{T}^{\text{merged}}$
Init: $\theta_{0}^{\text{merged}}\!\leftarrow\!\theta_{0}$;
$U_{0}^{(l)}\!\leftarrow\!\emptyset$
for task $t=1$ to $T$ do
      $\triangleright$ create low-rank experts (Eqs. 3 and 4)
       $f_{t}=\text{SVD}~{}\!\bigl{(}\,\underbrace{\mathcal{P}^{(t-1)}(\theta_{t}-\theta_{0})}_{\text{use }\theta_{1}-\theta_{0}\text{ when }t=1}\bigr{)}=B_{t}A_{t}$
      $\triangleright$ add expert & initialize gate (Eq. 5)
       $\theta_{t}^{\text{merged}}=\theta_{t-1}^{\text{merged}}+g_{t}\!\cdot\!f_{t}$
      $\{W_{t}^{(g)},b_{t}^{(g)},S^{(0)}\}\leftarrow 0$
      for $m=1$ to total steps do
            $X\leftarrow\text{batch}(\mathcal{D}^{\text{seed}}_{t})$
            $\{\nabla_{W_{t}^{(g)}}L,\nabla_{b_{t}^{(g)}}L\}\leftarrow\nabla L_{t}(X,\theta_{t}^{\text{merged}},\theta_{t})$
            if $t>1$ then
                  $\triangleright$ project gradient onto null-space (Eqs. 8-11)
                   $S^{(m)}=\beta S^{(m-1)}+(1-\beta)r$
                  $\nabla_{W_{t}^{(g)}}L\leftarrow\widetilde{P}_{t-1}\,\nabla_{W_{t}^{(g)}}L$
             end if
            $W_{t}^{(g)}\leftarrow W_{t}^{(g)}-\eta\,\nabla_{W_{t}^{(g)}}L$
            $b_{t}^{(g)}\leftarrow b_{t}^{(g)}-\eta\,\nabla_{b_{t}^{(g)}}L$
       end for
      $\triangleright$ update dominant subspaces
       $U_{t}\leftarrow\operatorname{orthonorm}\!\bigl{[}U_{t-1}\mid\tilde{U}_{t}]$
end for
Algorithm 1 Mingle Procedure.
Adaptive Null-Space Relaxation.  To restore plasticity, we replace the all-one eigenvalues of $P_{t-1}$ with soft coefficients learned online.
(i) Interference statistics.  For each column $u_{p}^{(l)}$ of $U_{t-1}^{(l)}$ we measure instantaneous alignment:
$r_{p}^{(l)}=\|(\nabla L_{t})^{\top}u_{p}^{(l)}\|_{2}\;/\;\|\nabla L_{t}\|_{2}.$ (8)
We maintain per-direction interference scores $S^{(m,l)}\in\mathbb{R}^{k}$ (initialized to 00) by applying an exponential moving average at each optimization step $m$:
$S^{(m,l)}=\beta\,S^{(m-1,l)}+(1-\beta)\,r^{(l)},$ (9)
which suppresses stochastic gradient noise while preserving the dominant interference directions.
(ii) Adaptive shrinkage.  Each direction is attenuated by $\lambda^{(m,l)}=\exp(-\gamma\,S^{(m,l)})$ ($\gamma>0,\ \lambda^{(m,l)}\in(0,1]$). Let $\Lambda^{(m,l)}_{t-1}=\mathrm{diag}(\lambda_{1}^{(m,l)},\dots,\lambda_{(t-1)\cdot k}^{(m,l)})$. The relaxed projector becomes:
$\widetilde{P}_{t-1}^{(m,l)}=U_{t-1}^{(m,l)}\,\Lambda^{(m,l)}_{t-1}\,U_{t-1}^{(m,l)\top},$ (10)
interpolating smoothly between no protection ($\Lambda=0$) and the hard null projector ($\Lambda=I$).
(iii) Update rule.  We finally update:
$W_{t}^{(g,l)}\;\leftarrow\;W_{t}^{(g,l)}-\eta\,\widetilde{P}_{t-1}^{(l-1)}\,\nabla L_{t}^{(l)}.$ (11)
Relaxing the projector inevitably allows more residual interference than the hard null-space variant, yet empirically the increase is minor and is offset by markedly higher plasticity (Tab. 4), indicating a favorable stability–plasticity balance. The overall procedure is outlined in Algo. 1.
4 Experiments
We describe the experimental setup in Sec. 4.1, followed by the main results in Sec. 4.2 and further analysis and ablations in Sec. 4.3. Due to page limitations, detailed results are provided in the Appendix.
4.1 Experimental Setup
Datasets and Models. Following [68], we evaluate our method on a diverse set of image classification tasks using CLIP-ViT models [52] as the backbone. Three task groups comprising 8, 14, and 20 tasks are considered to assess scalability. Specifically, we use ViT-B/32, ViT-B/16, and ViT-L/14 models, each fine-tuned on up to 20 downstream tasks. All resulting model checkpoints are publicly available via FusionBench [65]. To evaluate robustness to task order, we repeat experiments over 10 random permutations using seeds ranging from 42 to 51. To enable comparison with conventional continual learning methods, we adopt the Multi-domain Task-Incremental Learning (MTIL) benchmark [93], where eleven tasks are presented in alphabetical order.
Implementation Details. We insert low-rank experts into the linear layers of the CLIP vision encoder. Two variants are considered: a full setup that modifies all attention and MLP linear layers, and a lightweight alternative applied only to attn.qkv and mlp.fc1. Unless stated otherwise, we adopt the full configuration. Each expert has rank $r=64$. The null-space constraint uses a subspace dimension $k=3$, with relaxation coefficients $\gamma=1$ and $\beta=0.99$. For each task, the merged model is adapted over 50 iterations using the Adam optimizer with a learning rate of 1e−4 and a batch size of 16. We draw 5 unlabeled test data per class, with no access to data from previous tasks.
Evaluation Metrics. We evaluate performance using average accuracy (ACC) and backward transfer (BWT) [36]. ACC measures the mean accuracy over all tasks using the final merged model: $\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}a_{i}(\theta_{t}^{\text{merged}}),$ where $a_{i}(\cdot)$ denotes the accuracy on task $i$. BWT quantifies the extent of forgetting by comparing performance on earlier tasks before and after merging: $\text{BWT}=\frac{1}{T-1}\sum_{i=1}^{T-1}\left[a_{i}(\theta_{t}^{\text{merged}})-a_{i}(\theta_{i}^{\text{merged}})\right].$
Table 1: Comparative results of continual merging methods, reporting average accuracy (ACC) and backward transfer (BWT) over ten task orders (mean±$\pm$±std). Best results are in bold; second-best are underlined. Mingle denotes a lightweight variant.
Method ViT-B/32 ViT-B/16 ViT-L/14
8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
Pre-Trained 48.1 56.9 55.6 55.4 62.0 59.8 64.9 69.1 65.6
Fine-Tuned 90.4 89.3 89.8 92.4 91.3 91.6 94.3 93.4 93.5
C. Fine-Tuned 79.8 67.4 62.6 82.9 72.2 68.2 90.0 70.9 77.7
ACC (%) ↑
Average (SWA) [25] 66.3 $\pm$0.0 65.4 $\pm$0.0 61.1 $\pm$0.0 72.3 $\pm$0.0 69.7 $\pm$0.0 64.8 $\pm$0.0 80.0 $\pm$0.0 77.5 $\pm$0.0 71.1 $\pm$0.0
C. Task Arithmetic [23] 67.5 $\pm$0.0 66.5 $\pm$0.0 60.0 $\pm$0.0 77.1 $\pm$0.0 70.9 $\pm$0.6 64.2 $\pm$0.0 82.1 $\pm$0.0 77.9 $\pm$0.0 70.3 $\pm$0.0
C. Ties-Merging [82] 49.0 $\pm$10.2 66.2 $\pm$0.6 59.9 $\pm$0.7 66.8 $\pm$3.7 70.5 $\pm$0.8 63.0 $\pm$1.6 64.3 $\pm$7.0 78.0 $\pm$0.6 68.3 $\pm$0.9
MagMax-Ind [39] 70.7 $\pm$0.0 67.0 $\pm$0.0 61.2 $\pm$0.0 76.7 $\pm$1.8 67.0 $\pm$0.0 62.5 $\pm$0.0 83.4 $\pm$0.0 71.2 $\pm$0.0 71.2 $\pm$0.0
OPCM [68] 75.5 $\pm$0.5 71.9 $\pm$0.3 65.7 $\pm$0.2 81.8 $\pm$0.3 77.1 $\pm$0.5 70.3 $\pm$0.2 87.0 $\pm$0.4 83.5 $\pm$0.2 76.0 $\pm$0.2
Mingle 85.8 $\pm$0.8 81.6 $\pm$1.4 77.1 $\pm$2.0 88.3 $\pm$0.6 84.9 $\pm$0.8 81.9 $\pm$0.9 - - -
Mingle 85.0 $\pm$0.5 81.7 $\pm$1.0 77.1 $\pm$1.3 87.0 $\pm$0.6 84.7 $\pm$1.0 81.6 $\pm$1.3 91.4 $\pm$0.3 89.2 $\pm$0.1 83.6 $\pm$0.6
BWT (%) ↑
Average (SWA) [25] -11.5 $\pm$2.2 -8.0 $\pm$1.3 -7.1 $\pm$2.1 -9.7 $\pm$1.5 -7.1 $\pm$1.4 -7.3 $\pm$1.7 -7.3 $\pm$1.4 -5.8 $\pm$1.0 -6.4 $\pm$1.5
C. Task Arithmetic [23] -9.6 $\pm$1.5 -1.3 $\pm$1.6 -3.4 $\pm$1.0 -4.2 $\pm$1.0 -1.3 $\pm$0.4 -3.6 $\pm$0.4 -7.1 $\pm$0.8 -1.8 $\pm$0.3 -3.3 $\pm$0.3
C. Ties-Merging [82] -15.3 $\pm$8.0 1.9 $\pm$0.6 -1.5 $\pm$0.7 -5.5 $\pm$0.4 1.4 $\pm$0.7 -1.5 $\pm$1.2 -13.0 $\pm$5.7 -1.1 $\pm$0.4 -2.9 $\pm$1.0
MagMax-Ind [39] -8.3 $\pm$1.3 -7.4 $\pm$1.4 -7.2 $\pm$1.6 -6.1 $\pm$1.3 -7.4 $\pm$2.0 -8.0 $\pm$2.2 -5.0 $\pm$0.8 -6.0 $\pm$2.1 -6.5 $\pm$2.1
OPCM [68] -6.3 $\pm$1.1 -6.0 $\pm$1.0 -7.8 $\pm$1.5 -4.8 $\pm$0.7 -5.1 $\pm$1.4 -6.3 $\pm$2.2 -2.6 $\pm$1.0 -4.3 $\pm$0.7 -6.5 $\pm$1.8
Mingle -0.6 $\pm$0.4 -1.1 $\pm$0.3 -2.2 $\pm$0.8 -0.4 $\pm$0.1 -0.9 $\pm$0.1 -1.9 $\pm$0.4 - - -
Mingle -0.1 $\pm$0.1 -0.4 $\pm$0.1 -1.3 $\pm$0.6 -0.1 $\pm$0.1 -0.3 $\pm$0.1 -1.0 $\pm$0.4 -0.2 $\pm$0.0 -0.4 $\pm$0.2 -1.5 $\pm$0.6
Table 2: Robustness results for ViT-B/32 continual merged on four tasks.
Method Clean Motion Impulse Gaussian Pixelate Spatter Contrast JPEG Avg.
ACC(%)
C. Layer-wise AdaMerging [83] 56.0 $\pm$5.3 47.5 $\pm$4.4 43.1 $\pm$2.3 43.3 $\pm$3.4 18.1 $\pm$4.7 46.6 $\pm$3.0 48.9 $\pm$4.8 49.1 $\pm$4.0 44.9
C. WEMOE [66] 3.4 $\pm$0.8 3.1 $\pm$0.4 4.3 $\pm$1.4 3.4 $\pm$1.4 3.0 $\pm$1.6 4.0 $\pm$0.9 3.3 $\pm$0.7 4.0 $\pm$1.2 3.6
C. Task Arithmetic [23] 77.5 $\pm$0.0 66.0 $\pm$0.0 58.9 $\pm$0.0 59.6 $\pm$0.0 29.7 $\pm$0.0 63.5 $\pm$0.0 66.0 $\pm$0.0 67.8 $\pm$0.0 61.1
MagMax-Ind [39] 79.1 $\pm$0.0 69.0 $\pm$0.0 60.6 $\pm$0.0 61.5 $\pm$0.0 33.0 $\pm$0.0 66.4 $\pm$0.0 68.6 $\pm$0.0 69.9 $\pm$0.0 63.5
OPCM [68] 83.6 $\pm$0.5 72.5 $\pm$0.6 64.7 $\pm$1.2 65.2 $\pm$1.2 35.2 $\pm$0.6 70.5 $\pm$0.5 72.5 $\pm$0.6 74.4 $\pm$0.3 67.3
Mingle 89.9 $\pm$0.4 82.8 $\pm$0.8 67.5 $\pm$2.0 70.7 $\pm$1.2 37.9 $\pm$0.4 77.0 $\pm$0.7 80.1 $\pm$0.8 82.9 $\pm$0.9 73.2
BWT(%)
C. Layer-wise AdaMerging [83] -38.0 $\pm$7.1 -37.3 $\pm$5.9 -22.2 $\pm$3.0 -25.2 $\pm$4.5 -20.8 $\pm$6.3 -28.6 $\pm$4.0 -34.7 $\pm$6.5 -36.1 $\pm$5.3 -29.5
C. WEMOE [66] -30.7 $\pm$3.1 -28.7 $\pm$3.8 -22.1 $\pm$11.6 -25.5 $\pm$9.8 -8.0 $\pm$4.5 -23.4 $\pm$11.0 -27.6 $\pm$5.6 -28.6 $\pm$5.2 -24.3
C. Task Arithmetic [23] -4.8 $\pm$0.9 -6.1 $\pm$1.2 -1.6 $\pm$3.0 -1.6 $\pm$1.7 -2.7 $\pm$1.5 -3.1 $\pm$2.5 -6.1 $\pm$1.2 -5.1 $\pm$0.8 -3.9
MagMax-Ind [39] -7.7 $\pm$0.8 -8.1 $\pm$1.5 -6.1 $\pm$4.9 -5.1 $\pm$3.7 -3.5 $\pm$3.0 -7.3 $\pm$2.9 -8.4 $\pm$1.6 -8.2 $\pm$1.2 -6.8
OPCM [68] -4.3 $\pm$1.8 -4.5 $\pm$2.8 -6.4 $\pm$7.1 -6.1 $\pm$4.3 -2.9 $\pm$0.9 -6.3 $\pm$2.9 -4.5 $\pm$2.8 -5.7 $\pm$1.5 -5.1
Mingle -0.2 $\pm$0.2 -0.1 $\pm$0.4 0.7 $\pm$1.0 0.6 $\pm$1.1 -0.2 $\pm$1.1 0.2 $\pm$0.7 0.0 $\pm$0.5 0.5 $\pm$0.5 -0.2
4.2 Main Results
Overall Performance  As shown in Tab. 1, Mingle substantially outperforms existing continual merging approaches across all CLIP backbones and task counts, achieving the highest accuracy while nearly eliminating backward forgetting. These results confirm that our method consistently delivers both strong task accuracy and long-term stability under diverse continual scenarios.
Table 3: Comparison of last accuracy (%) with conventional CL approaches on MTIL benchmark.
Method
Aircraft
Caltech101
CIFAR100
DTD
EuroSAT
Flowers
Food101
MNIST
Pets
Cars
SUN397
 Avg.
Conventional CL (Sequential fine-tuned)
WiSE-FT [77] 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
ZSCL [93] 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
MoE-Adapter [85] 49.8 92.2 86.1 78.1 95.7 94.3 89.5 98.1 89.9 81.6 80.0 85.0
DIKI [69] 45.2 95.7 86.3 72.9 98.0 97.0 89.2 99.4 94.2 81.6 76.6 85.1
AwoForget [92] 42.4 92.7 83.2 73.2 97.0 91.8 92.2 99.1 93.9 87.4 82.6 85.0
Dual-RAIL [81] 52.5 96.8 83.3 80.1 96.4 99.0 89.9 98.8 93.5 85.5 79.2 86.8
MagMax [39] 40.2 96.1 81.1 72.0 97.8 76.3 88.4 99.2 93.0 70.5 68.9 80.3
Mingle-Seq 58.7 97.5 87.2 79.7 97.3 87.2 90.1 99.6 93.0 80.4 73.3 85.8
Continual Merging (Independent fine-tuned)
Average (SWA) [25] 26.5 92.3 74.3 48.4 73.7 74.0 87.1 84.0 91.2 67.5 68.5 71.6
C. TA [23] 26.6 92.5 74.5 48.7 74.3 74.4 87.0 85.5 91.2 67.7 68.6 71.9
C. Ties [82] 30.5 94.0 74.8 49.8 71.7 73.8 87.3 81.5 90.6 67.0 67.9 71.7
MagMax-Ind [39] 29.9 93.7 78.4 46.1 58.3 68.1 86.8 82.8 91.4 62.7 69.3 69.8
OPCM [68] 35.7 95.9 77.0 54.6 90.3 76.4 87.1 96.3 93.3 70.1 70.5 77.0
Mingle 54.2 97.3 79.7 72.3 96.0 86.7 88.7 99.3 93.9 73.1 71.6 83.0
Table 4: Ablation study of Mingle with CLIP ViT-B/16 over 8, 14, and 20 tasks.
Test-Time Frozen Null-Space Adaptive ACC(%) BWT(%)
Adaptation Old Gate Constrained Gate Relaxation 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
78.7 $\pm$0.1 76.4 $\pm$1.0 70.6 $\pm$0.4 -0.5 $\pm$0.1 -1.0 $\pm$0.3 -1.3 $\pm$0.3
86.4 $\pm$5.3 81.7 $\pm$2.3 76.7 $\pm$1.3 -6.0 $\pm$5.2 -7.7 $\pm$-3.4 -12.8 $\pm$1.3
87.4 $\pm$0.4 81.3 $\pm$0.8 76.2 $\pm$1.3 -2.3 $\pm$0.5 -4.3 $\pm$0.8 -6.8 $\pm$0.9
86.0 $\pm$1.5 83.5 $\pm$0.9 78.3 $\pm$1.7 -0.1 $\pm$0.1 -0.1 $\pm$0.1 -0.2 $\pm$0.1
88.3 $\pm$0.6 84.9 $\pm$0.8 81.9 $\pm$0.9 -0.4 $\pm$0.1 -0.9 $\pm$0.1 -1.9 $\pm$0.4
Robustness to Test-Time Distribution Shifts. Following prior work [83, 66], we evaluate Mingle on seven corruptions (motion blur, impulse noise, Gaussian noise, pixelate, spatter, contrast, JPEG) and report results in Tab. 2. It preserves high accuracy and near-zero or even positive BWT, outperforming all baselines, whereas direct application of SOTA TTA-based merging (WEMOE, AdaMerging) in a continual setting fails without tailored designs to continual setup.
Figure 3: Efficiency analysis on TTA time, trainable/full parameter count, and accuracy for CLIP-ViT-B/32 merged on eight tasks.
Configuration Time T. Param F. Param ACC
attn & mlp ($r=64$) 78 s 83.0 K 173.1 M 85.8
attn.qkv_proj ($r=64$) 61 s 27.7 K 116.0 M 69.9
attn.out_proj ($r=64$) 47 s 9.3 K 97.0 M 53.9
mlp.fc1 ($r=64$) 48 s 9.0 K 111.1 M 82.6
mlp.fc2 ($r=64$) 48 s 36.8 K 111.3 M 70.1
qkv & fc1 ($r=32$) 65 s 36.9 K 113.7 M 84.5
qkv & fc1 ($r=64$) 65 s 36.9 K 139.7 M 85.0
qkv & fc1 ($r=128$) 65 s 36.9 K 191.6 M 85.1
qkv & fc1 ($r=768$) 70 s 36.9 K 710.3 M 83.7
Comparison with Conventional CL. Tab. 3 evaluates two CL paradigms on the MTIL benchmark: conventional CL, where each task model is fine-tuned from its immediate predecessor; and continual merging, which fine-tunes each task model independently before fusion, eliminating inter-model dependencies and enabling flexible task ordering and model reuse. Within the merging family, Mingle sets a new state-of-the-art, and when integrated into a sequential fine-tuning pipeline, it matches the performance of SOTA CL methods. This demonstrates both its strength as a fusion strategy and its versatility across different training regimes.
4.3 Ablation Results and Analysis
Ablation Study.
We explore the contribution of each component in Tab. 4. Row 1 shows a fixed-weight merging of low-rank experts as our baseline. In Row 2, adding TTA boosts ACC substantially but at the cost of worsening BWT with more tasks. Row 3 demonstrates that freezing earlier gates curbs forgetting while retaining ACC gains. Row 4 then applies null-space constraints, yielding further BWT improvements. Finally, Row 5 presents the full method with adaptive relaxation, which best harmonizes accuracy and long-term stability.
Computation and Parameter Efficiency.  Tab. 3 compares different expert insertion layers and rank settings for CLIP-ViT-B/32 on eight tasks. Embedding experts in both attn and mlp layers achieves the highest accuracy at the cost of increased TTA time and parameter overhead. A hybrid scheme (qkv & fc1, $r=64$) delivers a favorable balance. In the rank sweep, moving from $r=32$ to $64$ yields a clear ACC gain, while further increasing $r$ leads to diminishing or negative returns. Experiments were run on a single RTX 4090 GPU. These results highlight Mingle’s flexibility under diverse resource budgets.
Number of Seed Samples  The quantity of seed samples plays a central role in determining the reliability and efficiency of TTA. We examine how varying the number of seed samples per class impacts performance with CLIP ViT-B/16. As shown in Fig. 6, even a single sample per class yields strong results, with only minor gains from additional samples. This indicates that TTA is largely robust to sample quantity. We also observe reduced variance across task orders as more samples are used, motivating our choice of five samples per class as a balanced trade-off between performance and efficiency.
Figure 4: Gate activations across eight tasks under varying $\gamma$. Each subplot shows one gate; curves and shaded areas indicate mean and std across layers. Gray bars mark the gate’s training task. Lower $\gamma$ leads to stronger suppression on prior tasks.
(a) Vary $\gamma$, $\beta=0.99$, $k=3$
(b) Vary $\beta$, $\gamma=1$, $k=3$
(c) Vary $k$, $\gamma=0$
Figure 5: Sensitivity analysis of the null-space constrained gating w.r.t. hyper-parameters $\beta$, $\gamma$, and $k$.
Figure 6: Analysis of seed sample quantity on performance.
Hyper-parameter of Gate.  We study the effect of three key hyper-parameters: $\gamma$, $\beta$, and $k$. $\gamma$ controls the strength of null-space suppression; smaller values lead to stronger attenuation of activations on prior tasks, reducing forgetting (Fig. 4, 5(a)). As shown in Fig. 5(b), $\beta$ regulates the smoothness of the EMA used to accumulate interference signals. A moderate setting ($\beta=0.99$) balances responsiveness and stability. Smaller $\beta$ amplifies noise sensitivity, while larger $\beta$ slows detection of interference. $k$ determines the number of principal directions retained per task; performance saturates at $k=3$ (Fig. 5(c)), indicating that a small number of task-specific directions suffices.
5 Conclusions
We proposed Mingle, a test-time continual model merging (TTCMM) approach that integrates test-time adaptation into continual model merging via a mixture-of-experts architecture and Adaptive Null-Space Constrained Gating. Extensive empirical evaluations confirmed that Mingle significantly improves generalization performance and mitigates catastrophic forgetting, consistently surpassing previous state-of-the-art methods. These results highlight Mingle’s capability to efficiently adapt to new tasks without revisiting original training data, demonstrating the practical advantages of our proposed TTCMM framework in realistic continual learning applications.
Appendix
The appendix is organized into sections, each providing additional information and details.
A  Theoretical Risk Comparison: Dynamic MoE vs. Static Averaging 16
B  Additional Descriptions 18
    B.1  Details of Dataset and Task Settings 18
    B.2  Details of Downstream Models 19
    B.3  Details of Baselines 19
C  Additional Results 21
    C.1  Detailed Overall Performance Results 21
    C.2  Accuracy Trends Across Sequential Tasks 21
    C.3  Detailed Results Under Distribution Shifts 25
    C.4  More Visualizations of Gate Activations and Relaxation Effect 25
D  Discussions 27
    D.1  Limitations 27
    D.2  Broader Impacts 28
Appendix A Theoretical Risk Comparison: Dynamic MoE vs. Static Averaging
Problem Setup and Definitions
Consider $T$ independent tasks, each associated with a data distribution $D_{t}$ for $t=1,\dots,T$. For each task $t$, a pre-trained expert model $f_{t}(x)$ outputs a probability distribution over classes, trained specifically on $D_{t}$. The overall data distribution $D$ is a mixture of these tasks, where an example $(x,y)$ is drawn from task $t$ with prior probability $P(t)$, and then $(x,y)\sim D_{t}$. The expected risk of a predictive model $h(x)$ is defined as:
$R(h)=\mathbb{E}_{(x,y)\sim D}\big{[}\ell(h(x),y)\big{]}=\sum_{t=1}^{T}P(t)\,\mathbb{E}_{(x,y)\sim D_{t}}\big{[}\ell(h(x),y)\big{]},$ (12)
where $\ell(h(x),y)$ is a loss function (e.g., cross-entropy or 0-1 loss).
We compare two methods to combine the experts into a final prediction $h(x)$:
  • Static Averaging: Defined as $h_{\text{static}}(x)=\sum_{i=1}^{T}\alpha_{i}f_{i}(x)$, where $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{T})$ is a fixed weight vector independent of $x$, typically with $\alpha_{i}\geq 0$ and $\sum_{i}\alpha_{i}=1$ for probability outputs.
  • Dynamic Mixture-of-Experts (MoE): Defined as $h_{\text{MoE}}(x)=f_{i^{*}(x)}(x)$, where $i^{*}(x)=\arg\max_{i}g_{i}(x)$ and $g(x)=(g_{1}(x),\dots,g_{T}(x))$ is a gating function that selects one expert per input (hard routing). The gating is subject to routing noise, modeled below.
Routing Noise Model
For each input $x$ drawn from $D_{t}$, the true task is $t$, and the ideal expert is $f_{t}$. The gating selects the correct expert $i^{*}(x)=t$ with probability $1-\varepsilon_{t}$, and an incorrect expert $i^{*}(x)\neq t$ with probability $\varepsilon_{t}=P(i^{*}(x)\neq t\mid x\sim D_{t})$, the task-specific routing error rate. On error, the gating selects a random expert from $\{1,\dots,T\}\setminus\{t\}$ uniformly. Define:
  • $R_{t}(i)=\mathbb{E}_{(x,y)\sim D_{t}}[\ell(f_{i}(x),y)]$, the risk of expert $i$ on task $t$.
  • $R_{\text{ideal}}=\sum_{t=1}^{T}P(t)R_{t}(t)$, the risk with perfect routing.
  • $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$, the average risk of incorrect experts on task $t$.
  • $\varepsilon=\sum_{t=1}^{T}P(t)\varepsilon_{t}$, the overall routing error rate.
Theorem 2 (Dynamic MoE versus Static Averaging).
Let $\{(D_{t},f_{t})\}_{t=1}^{T}$ be $T$ independent tasks with priors $P(t)$ and per-task risks $R_{t}(i)$. For any static mixture $h_{\mathrm{static}}(x)=\sum_{i=1}^{T}\alpha_{i}\,f_{i}(x)$ and any hard-routed MoE $h_{\mathrm{MoE}}(x)=f_{i^{\star}(x)}(x)$ with task-specific routing errors $\varepsilon_{t}$:
$R(h_{\mathrm{MoE}})\;=\;R_{\mathrm{ideal}}+\sum_{t=1}^{T}P(t)\,\varepsilon_{t}\bigl{(}R_{\text{wrong},t}-R_{t}(t)\bigr{)},$ (13)
where $R_{\mathrm{ideal}}=\sum_{t}P(t)R_{t}(t)$ and $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$. Moreover,
  1. 1.
    (Perfect routing) If $\varepsilon_{t}=0$ for all $t$, then $\inf_{g}R(h_{\mathrm{MoE}})\;<\;\inf_{\boldsymbol{\alpha}}R(h_{\mathrm{static}})$ whenever at least two tasks disagree on their best expert.
  2. 2.
    (Noisy routing)  If $\sum_{t}P(t)\varepsilon_{t}\bigl{(}R_{\text{wrong},t}-R_{t}(t)\bigr{)}<R^{*}_{\mathrm{static}}-R_{\mathrm{ideal}},$ where $R^{*}_{\text{static}}=\inf_{\boldsymbol{\alpha}}R(h_{\text{static}})$, then the MoE still attains lower risk than any static mixture.
Proof.
The proof proceeds in three parts: (1) deriving the MoE risk with routing noise, (2) proving the optimal gating case, and (3) establishing the condition for MoE superiority under routing noise.
Step 1: MoE Risk with Routing Noise
The MoE prediction is $h_{\text{MoE}}(x)=f_{i^{*}(x)}(x)$, where $i^{*}(x)=\arg\max_{i}g_{i}(x)$. The expected risk is:
$R(h_{\text{MoE}})=\mathbb{E}_{(x,y)\sim D}[\ell(f_{i^{*}(x)}(x),y)]=\sum_{t=1}^{T}P(t)\,\mathbb{E}_{(x,y)\sim D_{t}}[\ell(f_{i^{*}(x)}(x),y)].$ (14)
For task $t$, condition on routing correctness:
  • Correct routing ($i^{*}(x)=t$): Probability $1-\varepsilon_{t}$, risk $R_{t}(t)$.
  • Incorrect routing ($i^{*}(x)\neq t$): Probability $\varepsilon_{t}$, selects a random expert from $\{1,\dots,T\}\setminus\{t\}$, with average risk $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$.
The expected risk on task $t$ is:
$\mathbb{E}_{(x,y)\sim D_{t}}[\ell(f_{i^{*}(x)}(x),y)]=(1-\varepsilon_{t})R_{t}(t)+\varepsilon_{t}R_{\text{wrong},t}.$ (15)
Thus, the total risk is:
$R(h_{\text{MoE}})=\sum_{t=1}^{T}P(t)\big{[}(1-\varepsilon_{t})R_{t}(t)+\varepsilon_{t}R_{\text{wrong},t}\big{]}.$ (16)
Rewrite:
$R(h_{\text{MoE}})=\sum_{t=1}^{T}P(t)R_{t}(t)+\sum_{t=1}^{T}P(t)\varepsilon_{t}(R_{\text{wrong},t}-R_{t}(t))=R_{\text{ideal}}+\sum_{t=1}^{T}P(t)\varepsilon_{t}\delta_{t},$ (17)
where $\delta_{t}=R_{\text{wrong},t}-R_{t}(t)>0$ is the risk increase due to misrouting on task $t$.
Step 2: Optimal Gating (No Routing Noise)
Assume an oracle gating function with $\varepsilon_{t}=0$ for all $t$, so $i^{*}(x)=t$ for all $x\sim D_{t}$. Then:
$R(h_{\text{MoE}})=\sum_{t=1}^{T}P(t)R_{t}(t)=R_{\text{ideal}}.$ (18)
Define hypothesis classes:
$\mathcal{H}_{\text{static}}=\left\{x\mapsto\sum_{i=1}^{T}\alpha_{i}f_{i}(x)\mid\boldsymbol{\alpha}\in\mathbb{R}^{T}\right\},$ (19)
$\mathcal{H}_{\text{MoE}}=\left\{x\mapsto f_{i^{*}(x)}(x)\mid i^{*}(x)=\arg\max_{i}g_{i}(x),g:\mathcal{X}\to\mathbb{R}^{T}\right\}.$ (20)
Any static model $h_{\text{static}}(x)=\sum_{i=1}^{T}\alpha_{i}f_{i}(x)$ can be approximated by an MoE with $g(x)$ assigning constant weights, so $\mathcal{H}_{\text{static}}\subseteq\mathcal{H}_{\text{MoE}}$. Thus:
$R^{*}_{\text{MoE}}=\inf_{g(\cdot)}R(h_{\text{MoE}})\leq\inf_{\boldsymbol{\alpha}}R(h_{\text{static}})=R^{*}_{\text{static}}.$ (21)
Under task heterogeneity ($R_{t}(t)<R_{t}(s)$ and $R_{s}(s)<R_{s}(t)$ for some $t\neq s$), the ideal MoE routes each $x\sim D_{t}$ to $f_{t}$, achieving:
$R_{\text{ideal}}=\sum_{t=1}^{T}P(t)R_{t}(t).$ (22)
For static averaging:
$R(h_{\text{static}})=\sum_{t=1}^{T}P(t)\mathbb{E}_{(x,y)\sim D_{t}}\left[\ell\left(\sum_{i=1}^{T}\alpha_{i}f_{i}(x),y\right)\right].$ (23)
Since $\ell$ is convex (e.g., cross-entropy), Jensen’s inequality implies:
$\mathbb{E}_{D_{t}}\left[\ell\left(\sum_{i}\alpha_{i}f_{i}(x),y\right)\right]\geq R_{t}(t),$ (24)
with strict inequality unless $\alpha_{t}=1$ and $\alpha_{i}=0$ for $i\neq t$, which cannot hold for all tasks simultaneously under heterogeneity. Thus:
$R^{*}_{\text{static}}>R_{\text{ideal}}=R^{*}_{\text{MoE}}.$ (25)
Step 3: MoE Superiority with Routing Noise
Let $\gamma=R^{*}_{\text{static}}-R_{\text{ideal}}>0$ under task heterogeneity. The MoE outperforms the static model if:
$R(h_{\text{MoE}})<R^{*}_{\text{static}}.$ (26)
Substitute:
$R_{\text{ideal}}+\sum_{t=1}^{T}P(t)\varepsilon_{t}\delta_{t}<R_{\text{ideal}}+\gamma.$ (27)
Thus:
$\sum_{t=1}^{T}P(t)\varepsilon_{t}\delta_{t}<\gamma=R^{*}_{\text{static}}-R_{\text{ideal}}.$ (28)
Since $\delta_{t}=R_{\text{wrong},t}-R_{t}(t)$, the condition is:
$\sum_{t=1}^{T}P(t)\varepsilon_{t}(R_{\text{wrong},t}-R_{t}(t))<R^{*}_{\text{static}}-R_{\text{ideal}}.$ (29)
If this holds, the MoE’s risk, despite routing noise, remains below the best static risk. ∎
Conclusion: The MoE risk is $R_{\text{ideal}}+\sum_{t=1}^{T}P(t)\varepsilon_{t}(R_{\text{wrong},t}-R_{t}(t))$, and it outperforms static averaging when routing noise is sufficiently small relative to the static model’s suboptimality. The optimal gating case confirms $R^{*}_{\text{MoE}}\leq R^{*}_{\text{static}}$, with strict inequality under task heterogeneity.
Appendix B Additional Descriptions
B.1 Details of Dataset and Task Settings
Dataset Details
Following prior works [68], we evaluate continual model merging on twenty publicly available image classification datasets, including SUN397 [80], Stanford Cars [31], RESISC45 [4], EuroSAT [18], SVHN [44], GTSRB [62], MNIST [33], DTD [6], Flowers102 [45], PCAM [71], FER2013 [17], Oxford-IIIT Pet [48], STL-10 [8], CIFAR-100 and CIFAR-10 [32], Food-101 [3], Fashion-MNIST [79], EMNIST [9], KMNIST [7], and Rendered SST-2 [60].
Table 5: Extended downstream datasets used in our experiments.
Dataset #Classes #Train (k) #Test (k) Task
SUN397 287 19.9 19.9 Scene category
Stanford Cars 196 8.1 8.0 Car series
RESISC45 45 18.9 6.3 Remote–sensing scene
EuroSAT 10 21.6 2.7 Satellite land-use
SVHN 10 73.3 26.0 Digit recognition
GTSRB 43 39.2 12.6 Traffic sign
MNIST 10 60 10 Hand-written digit
DTD 47 3.8 1.9 Texture recognition
Flowers102 102 1.0 6.1 Flower species
PCAM 2 262 32.8 Tumour detection
FER2013 7 28.7 3.6 Facial emotion
Oxford IIIT Pet 37 3.7 3.7 Animal species
STL10 10 5 8 Object recognition
CIFAR-100 100 50 10 Natural object
CIFAR-10 10 50 10 Natural object
Food101 101 75.8 25.3 Food type
Fashion-MNIST 10 60 10 Fashion product
EMNIST (digits) 10 60 10 Hand-written digit
KMNIST 10 60 10 Kuzushiji character
Rendered SST-2 2 6.9 1.8 Rendered sentiment
Task Grouping
We group the 20 datasets into three progressive task sets and evaluate the merged models using average accuracy (ACC) and backward transfer (BWT) metrics. For each task group, we perform 10 experiments using different task sequences (listed in Tab.  6), and report both the mean and standard deviation of the results to ensure robustness and consistency.
  • 8-task group: (1) SUN397, (2) Stanford Cars, (3) RESISC45, (4) EuroSAT, (5) SVHN, (6) GTSRB, (7) MNIST, (8) DTD.
  • 14-task group: (1) SUN397, (2) Stanford Cars, (3) RESISC45, (4) EuroSAT, (5) SVHN, (6) GTSRB, (7) MNIST, (8) DTD, (9) Flowers102, (10) PCAM, (11) FER2013, (12) OxfordIIITPet, (13) STL10, (14) CIFAR100.
  • 20-task group: (1) SUN397, (2) Stanford Cars, (3) RESISC45, (4) EuroSAT, (5) SVHN, (6) GTSRB, (7) MNIST, (8) DTD, (9) Flowers102, (10) PCAM, (11) FER2013, (12) OxfordIIITPet, (13) STL10, (14) CIFAR100, (15) CIFAR10, (16) Food101, (17) FashionMNIST, (18) EMNIST, (19) KMNIST, (20) RenderedSST2.
Table 6: Dataset orderings used for experiments in each task group.
Group Order Dataset Order (by ID)
[…]
B.2 Details of Downstream Models
In this section, we present the evaluation setup for pre-trained and fine-tuned models. As shown in Tab. 7, we evaluate the zero-shot accuracy of the original CLIP-ViT models and the performance of fine-tuned models on the test sets of various downstream tasks. The fine-tuned checkpoints are obtained directly from Hugging Face (https://huggingface.co/tanganke), where each model has been fine-tuned on task-specific training data using a standard protocol. The visual encoder is updated during fine-tuning, while the classification head is fixed and initialized from the pre-trained text encoder. The fine-tuning setup follows a standard configuration: cross-entropy loss, Adam optimizer, cosine annealing learning rate schedule with a peak learning rate of 1e-5, batch size 128, and 4000 training steps.
Table 7: Test set accuracy of the pre-trained model and individual fine-tuned models on different downstream tasks.
Model
SUN397
Cars
RESISC45
EuroSAT
SVHN
GTSRB
MNIST
DTD
Flowers102
PCAM
CLIP-ViT-B/32
Pre-trained 63.2 59.6 60.3 45.0 31.6 32.5 48.3 44.2 66.4 60.6
Fine-tuned 74.9 78.5 95.1 99.1 97.3 98.9 99.6 79.7 88.6 88.0
CLIP-ViT-B/16
Pre-trained 65.5 64.7 66.4 54.1 52.0 43.5 51.7 45.0 71.3 54.0
Fine-tuned 78.9 85.9 96.6 99.0 97.6 99.0 99.7 82.3 94.9 90.6
CLIP-ViT-L/14
Pre-trained 68.2 77.9 71.3 61.2 58.4 50.5 76.3 55.5 79.2 51.2
Fine-tuned 82.8 92.8 97.4 99.1 97.9 99.2 99.8 85.5 97.7 91.1
Model
FER2013
OxfordIIITPet
STL10
CIFAR100
CIFAR10
Food101
FashionMNIST
EMNIST
KMNIST
RenderedSST2
CLIP-ViT-B/32
Pre-trained 41.3 83.3 97.1 63.7 89.8 82.4 63.0 12.0 10.0 58.6
Fine-tuned 71.6 92.5 97.5 88.4 97.6 88.4 94.7 95.6 98.2 71.3
CLIP-ViT-B/16
Pre-trained 46.4 88.4 98.3 66.3 90.8 87.0 67.3 12.4 11.2 60.6
Fine-tuned 72.8 94.5 98.2 88.8 98.3 91.9 94.5 95.3 98.1 75.7
CLIP-ViT-L/14
Pre-trained 50.0 93.2 99.4 75.1 95.6 91.2 67.0 12.3 9.7 68.9
Fine-tuned 75.9 95.7 99.2 93.0 99.1 94.8 95.3 95.4 98.3 80.5
B.3 Details of Baselines
Our experiments involve the following comparison methods and our method:
  • Stochastic Weight Averaging (SWA). A simple model averaging technique to stabilize optimization and improve generalization [25]. At each step $t$, the model parameters are averaged across previous checkpoint: $\theta^{\text{SWA}}_{t}=\frac{1}{t}\left[\theta^{\text{SWA}}_{t-1}(t-1)+\theta^{\text{SWA}}_{t}\right]$. This approach can be interpreted as a form of uniform model ensembling. While conceptually straightforward, SWA treats all checkpoints equally and does not account for inter-task conflicts.
  • Continual Task Arithmetic (C. TA). A training-free merging strategy that linearly combines task-specific fine-tuned models with a shared pre-trained model [23]. It computes the merged parameters as $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda(\theta_{t}-\theta_{0})$, where $\lambda$ is a scaling factor. TA is computationally efficient and easy to apply, but sensitive to $\lambda$ and prone to destructive interference when merging dissimilar tasks.
  • Continual Ties-Merging (C. Ties). An extension of Task Arithmetic that reduces parameter-level redundancy and sign conflicts during model merging [82]. For task $t$, the difference vector $\Delta\theta_{t}=\theta_{t}-\theta_{0}$ is trimmed and sign-normalized to obtain $\Delta\theta^{\text{Ties}}_{t}=\mathrm{Ties}{\left(\Delta\theta^{\text{Ties}}_{t-1},\Delta\theta_{t}\right)}$, and the merged model is given by $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda\Delta\theta^{\text{Ties}}_{t}$.
  • Orthogonal Projection-based Continual Merging (OPCM). A projection-based scheme to mitigate task interference by enforcing orthogonality between parameter updates [68]. Specifically, each $\Delta\theta_{t}$ is projected onto the orthogonal complement of the subspace spanned by previous updates: $\theta^{\text{merged}}_{t}=\theta_{0}+\frac{1}{\lambda_{t}}\left[\lambda_{t-1}\Delta\theta^{\text{merged}}_{t-1}+\mathcal{P}^{(t-1)}(\Delta\theta_{t})\right]$, where $\mathcal{P}^{(t-1)}$ denotes the orthogonal projection.
  • Maximum Magnitude Selection (MagMax). An extension of Task Arithmetic that, for each parameter dimension, selects the update with the larger absolute value: $\Delta\theta^{\text{MagMax}}_{t}=\mathrm{MagMax}{\left(\Delta\theta^{\text{MagMax}}_{t-1},\Delta\theta_{t}\right)}$, and the merged model is given by $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda\Delta\theta^{\text{MagMax}}_{t}$.
Appendix C Additional Results
In this section, we provide additional experimental results to support the findings reported in the main paper. Specifically, we include: (1) detailed overall performance results (C.1); (2) accuracy trends across sequential tasks (C.2); (3) detailed results under distribution shifts (C.3); and (4) extended visualizations of gate activations and hyperparameter effects (C.4).
C.1 Detailed Overall Performance Results
Tab. 8 expands on the average results in Tab. 1 by reporting per-task average accuracy after continually merging 20 tasks. We compare six methods, SWA, Task Arithmetic, Ties-Merging, MagMax-IND, OPCM, and our proposed Mingle across three CLIP-ViT backbones (B/32, B/16, L/14). Mingle achieves the highest accuracy on most tasks. These fine-grained results reinforce the main paper’s findings, highlighting Mingle’s ability to improve performance on continual model merging.
Table 8: Test set accuracy comparisons on different downstream tasks.
Model
SUN397
Cars
RESISC45
EuroSAT
SVHN
GTSRB
MNIST
DTD
Flowers102
PCAM
ViT-B/32
C. Fine-Tuned 53.9 38.2 64.7 98.7 45.4 34.4 86.7 58.4 57.5 67.7
Average (SWA) 64.2 59.6 64.8 60.9 47.3 43.1 71.8 46.4 66.5 63.9
C.TA 62.0 53.7 60.9 58.1 48.5 48.9 79.4 46.1 61.1 73.4
C.TIES 62.5 49.1 55.8 50.9 54.6 49.3 82.0 46.7 58.5 69.9
MagMax-Ind 63.6 53.1 59.7 49.1 53.8 53.1 79.8 43.2 56.9 75.1
OPCM 64.4 51.1 66.0 71.7 66.1 56.0 90.2 40.4 64.9 80.2
Mingle (Ours) 67.8 58.3 83.5 90.0 82.9 91.8 98.0 65.3 74.0 66.9
ViT-B/16
C. Fine-Tuned 62.7 58.0 67.6 99.1 46.0 29.2 93.9 61.9 64.1 75.2
Average (SWA) 67.1 64.6 69.3 63.4 62.4 52.7 80.7 46.6 71.8 63.1
C.TA 65.8 57.5 63.8 59.5 64.7 54.0 88.0 45.3 67.5 67.1
C.TIES 64.2 52.9 60.9 53.0 62.8 48.8 88.4 45.0 61.3 68.5
MagMax-Ind 65.8 51.8 57.8 42.6 54.4 43.7 83.0 42.8 60.4 69.8
OPCM 67.9 55.9 73.7 77.5 74.4 63.2 94.1 49.2 72.3 79.6
Mingle (Ours) 71.9 64.4 83.0 84.7 83.8 85.3 92.3 60.7 83.5 77.7
ViT-L/14
C. Fine-Tuned 69.5 73.6 78.3 99.2 59.3 49.3 98.6 69.7 83.2 78.3
Average (SWA) 70.7 77.7 76.4 75.3 69.5 62.1 93.7 57.7 80.0 73.6
C.TA 70.4 74.1 73.9 66.3 69.9 65.6 95.1 56.6 78.6 70.4
C.TIES 69.7 70.3 65.3 47.9 76.1 63.6 94.7 54.4 77.9 72.3
MagMax-Ind 73.1 73.7 75.6 64.6 73.7 68.8 94.6 56.1 78.0 71.7
OPCM 73.1 78.3 82.4 80.2 80.8 80.4 97.4 61.6 84.8 76.3
Mingle (Ours) 74.5 85.9 90.5 92.5 90.1 92.7 98.1 69.2 95.7 74.0
Model
FER2013
OxfordIIITPet
STL10
CIFAR100
CIFAR10
Food101
FashionMNIST
EMNIST
KMNIST
RenderedSST2
ViT-B/32
C. Fine-Tuned 58.3 68.5 86.7 40.2 70.5 50.0 90.7 72.4 54.5 54.5
Average (SWA) 50.2 84.1 97.0 69.8 92.7 80.4 71.3 15.0 11.5 61.8
C.TA 51.4 82.3 94.9 64.6 91.4 71.9 73.9 17.8 12.2 59.9
C.TIES 49.5 81.3 95.2 63.7 91.2 70.2 73.7 17.8 16.9 59.8
MagMax-Ind 56.5 79.9 94.6 68.7 91.9 73.8 74.3 18.3 15.4 63.9
OPCM 55.8 82.9 95.9 67.6 92.8 74.0 76.3 22.4 18.3 64.6
Mingle (Ours) 65.0 85.5 97.0 72.6 94.1 81.5 85.4 50.4 65.2 67.1
ViT-B/16
C. Fine-Tuned 60.5 84.5 90.5 38.8 73.6 61.9 89.7 83.3 51.5 72.8
Average (SWA) 50.9 89.6 98.0 72.9 94.2 85.9 73.3 15.6 12.4 62.5
C.TA 50.7 89.3 97.0 68.0 93.1 80.3 75.7 18.1 16.7 61.8
C.TIES 50.4 87.9 96.3 63.1 91.7 78.0 75.0 23.4 24.9 61.5
MagMax-Ind 57.7 88.8 97.5 71.5 94.4 81.3 77.2 24.5 25.0 59.4
OPCM 59.5 91.8 97.7 73.2 94.7 83.1 81.3 26.5 23.4 66.8
Mingle (Ours) 65.1 92.2 97.4 74.3 94.8 87.3 84.8 67.1 81.7 71.7
ViT-L/14
C. Fine-Tuned 68.0 92.1 94.5 60.5 85.7 74.8 93.1 89.0 59.2 78.8
Average (SWA) 52.7 94.2 99.2 81.7 97.0 90.7 77.4 16.1 10.4 66.1
C.TA 55.7 94.2 98.6 79.1 91.6 87.6 80.8 17.6 10.6 63.6
C.TIES 57.6 93.5 97.8 74.0 95.6 84.7 79.7 20.2 12.6 58.4
MagMax-Ind 52.9 93.9 98.7 82.1 97.3 89.5 81.6 19.2 11.1 68.4
OPCM 61.8 95.4 99.2 83.0 97.8 90.9 86.0 26.4 14.7 71.0
Mingle (Ours) 67.9 96.0 99.4 84.7 97.8 92.4 88.8 53.0 57.1 75.5
C.2 Accuracy Trends Across Sequential Tasks
Fig. 7 provides a detailed view of accuracy throughout the continual merging process across different settings, showing both the performance on the current task and on previously encountered ones. The progressive accuracy drop across columns illustrates the degree of forgetting over time. Notably, Mingle consistently alleviates this degradation, demonstrating markedly reduced forgetting across the full task sequence. Fig. 8 further compares the average accuracy curves of Mingle and baseline methods on previously seen tasks after each new model is merged, using the CLIP ViT-B/16 backbone. Results are averaged over 10 random task orderings. Mingle consistently achieves the highest performance throughout the merging process, with its accuracy curve clearly dominating those of competing methods. Moreover, the narrower standard deviation bands indicate that Mingle is more robust to the task orders.
(a) ViT-B/32 on 8 Tasks
(b) ViT-B/16 on 8 Tasks
(c) ViT-L/14 on 8 Tasks
(d) ViT-B/16 on 14 Tasks
(e) ViT-L/14 on 14 Tasks
(f) ViT-L/14 on 20 Tasks
Figure 7: Accuracy matrices of Mingle (ViT-B/32, ViT-B/16, and ViT-L/14) under different task settings.
(a) 8 tasks
(b) 14 tasks
(c) 20 tasks
Figure 8: Sequential test accuracy curves of Mingle and baselines (C.TA, C.Ties, MagMax, OPCM) under different task settings. Shaded regions indicate standard deviation across 10 task orders.
Table 9: Robustness results when merging ViT-B/32 models on four tasks.
Clean Test Set Corruption: Motion Blur
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 65.3 49.7 65.4 43.8 56.0 64.2 25.6 62.5 37.5 47.5
C. WEMOE 0.5 8.1 2.6 2.5 3.4 0.5 8.0 1.8 2.3 3.2
C. Task Arithmetic 64.6 90.4 80.2 74.8 77.5 62.3 59.4 78.5 63.3 65.9
MagMax-Ind 63.1 89.2 81.7 82.5 79.1 61.4 62.1 80.0 72.6 69.0
OPCM 65.7 92.3 85.7 90.5 83.6 62.8 62.5 83.7 82.2 72.8
Mingle 74.4 96.5 91.5 97.3 89.9 73.2 70.5 91.9 95.8 82.9
Corruption: Impulse Noise Corruption: Gaussian Noise
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 60.5 30.1 56.3 25.5 43.1 62.3 25.6 59.7 25.6 43.3
C. WEMOE 0.5 11.2 2.3 3.2 4.3 0.5 8.1 2.4 2.8 3.4
C. Task Arithmetic 59.9 57.7 72.9 45.0 58.9 61.8 51.4 75.1 50.1 59.6
MagMax-Ind 59.2 56.3 74.3 52.5 60.6 60.6 51.7 77.0 56.5 61.5
OPCM 61.1 57.1 78.5 62.0 64.7 63.0 52.4 80.7 64.9 65.2
Mingle 69.6 28.0 86.1 86.1 67.5 72.0 38.5 89.4 82.9 70.7
Corruption: Pixelate Corruption: Spatter
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 3.4 16.5 13.5 39.2 18.1 61.3 34.1 58.2 32.8 46.6
C. WEMOE 0.5 6.3 2.5 2.5 3.0 0.5 10.1 2.7 2.6 4.0
C. Task Arithmetic 2.5 31.7 19.1 65.6 29.7 61.2 63.1 72.7 57.0 63.5
MagMax-Ind 2.6 36.1 19.3 74.0 33.0 60.0 64.9 74.8 66.1 66.4
OPCM 2.1 34.3 19.5 84.9 35.2 61.5 64.7 78.8 76.8 70.5
Mingle 2.3 35.6 18.5 95.1 37.9 70.1 57.8 86.2 93.9 77.0
Corruption: Contrast Corruption: JPEG Compression
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 61.8 26.0 63.1 44.8 48.9 65.1 29.6 65.4 36.4 49.1
C. WEMOE 0.5 7.5 2.3 3.0 3.3 0.5 10.5 2.4 2.7 4.0
C. Task Arithmetic 62.5 55.2 75.3 70.8 66.0 64.1 66.2 80.0 61.0 67.8
MagMax-Ind 61.3 58.0 76.9 78.2 68.6 62.5 67.7 81.1 68.5 69.9
OPCM 63.8 57.5 81.3 87.4 72.5 65.0 68.0 85.4 79.3 74.4
Mingle 72.4 60.1 90.4 97.3 80.1 73.7 73.5 92.0 92.4 82.9
C.3 Detailed Results Under Distribution Shifts
Tab. 9 expands on Tab. 2 by reporting per-dataset accuracy under both clean test conditions and seven common corruption types: motion blur, impulse noise, Gaussian noise, pixelation, spatter, contrast shift, and JPEG compression. We evaluate six merging methods, across four downstream tasks: Cars, EuroSAT, RESISC45, and GTSRB. This detailed breakdown complements the average results in the main paper, providing a more comprehensive assessment of robustness under test-time distribution shifts.
C.4 More Visualizations of Gate Activations and Relaxation Effect
We provide an extended ablation study on gate hyper-parameters, including visualizations of gate activations under 14-task (Fig. 9) and 20-task (Fig. 10) configurations, complementing the 8-task results presented in the main paper. The visualizations demonstrate that the null-space constraint remains effective as the number of tasks increases, consistently suppressing gate responses to inputs from previously seen tasks and thereby mitigating forgetting.
Figure 9: Visualization of gate activations across 14 tasks under varying $\gamma$ values. Each subplot corresponds to a gate, with curves and shaded regions denoting the mean and standard deviation of activations across layers. Gray bars mark the training dataset for each gate. Smaller $\gamma$ values result in stronger suppression of activations on previously learned tasks.
Figure 10: Visualization of gate activations across 20 tasks under varying $\gamma$ values.
Appendix D Discussions
D.1 Limitations
As with many model merging methods, our approach assumes that all independently fine-tuned models originate from a shared pretrained initialization. The extent to which this assumption influences merging performance remains unclear and warrants further investigation. In addition, our current experiments focus on merging models with identical backbone architectures (e.g., CLIP ViT-B/16). Although our use of LoRA-based expert offers some structural uniformity, which could potentially accommodate heterogeneous backbones, we have not yet explored this setting. Extending our framework to support diverse initialization points or architectural variants remains an open direction for future work.
D.2 Broader Impacts
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.