### v3

# Mingle: Mixture of Null-Space Gated Low-Rank Experts for Test-Time Continual Model Merging

Abstract
Continual model merging integrates independently fine-tuned models sequentially without access to the original training data, offering a scalable and efficient solution for continual learning. However, existing methods face two critical challenges: parameter interference among tasks, which leads to catastrophic forgetting, and limited adaptability to evolving test distributions. To address these issues, we introduce the task of Test-Time Continual Model Merging (TTCMM), which leverages a small set of unlabeled test samples during inference to alleviate parameter conflicts and handle distribution shifts. We propose Mingle, a novel framework for TTCMM. Mingle employs a mixture-of-experts architecture with parameter-efficient, low-rank experts, which enhances adaptability to evolving test distributions while dynamically merging models to mitigate conflicts. To further reduce forgetting, we propose Null-Space Constrained Gating, which restricts gating updates to subspaces orthogonal to prior task representations, thereby suppressing activations on old tasks and preserving past knowledge. We further introduce an Adaptive Relaxation Strategy that adjusts constraint strength dynamically based on interference signals observed during test-time adaptation, striking a balance between stability and adaptability. Extensive experiments on standard continual merging benchmarks demonstrate that Mingle achieves robust generalization, significantly reduces forgetting, and consistently surpasses previous state-of-the-art methods by 7–9% on average across diverse task orders. Our code is available at: https://github.com/zihuanqiu/MINGLE
1 Introduction
Continual learning aims to incrementally adapt machine learning models to new tasks without forgetting previously learned knowledge, addressing the critical challenge of catastrophic forgetting [43]. However, conventional continual learning approaches typically require continuous access to original training data, raising significant concerns about privacy and substantial computational overhead due to retraining efforts, thus limiting their applicability in dynamic, data-sensitive environments.
To address these limitations, recent works have explored an alternative paradigm known as continual model merging (CMM), which sequentially integrates independently fine-tuned models directly in parameter space, without revisiting any training data [38, 51, 70]. CMM typically operates under a "merge-to-transfer" paradigm: given a pretrained model $\theta_{0}$ and independently fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$, a unified model is constructed sequentially by combining task-specific weight updates $\Delta\theta_{t}=\theta_{t}-\theta_{0}$ via weighted averaging or projection-based strategies [24, 81, 69].
Despite its advantages in scalability, data privacy, and distributed training capabilities [44, 14, 57], existing CMM methods still encounter critical issues, notably severe parameter interference between tasks and limited adaptability to evolving test distributions. This parameter interference arises because, as fine-tuned models are incrementally merged, overlapping or conflicting parameter updates accumulate, resulting in severe forgetting of previously learned tasks. To mitigate this interference, recent methods introduce structural constraints such as orthogonal projection [70, 78], model linearization [38, 69], and pruning-based sparsification [86, 90]. However, their effectiveness diminishes as task count grows and interference becomes increasingly entangled. Moreover, models merged across tasks often fail to generalize effectively, particularly when facing unseen or shifting task conditions. These flaws result in severe forgetting of earlier tasks and substantial performance gaps compared to the upper bound achieved by individually fine-tuned models. As shown in Fig. 1, TA [24] suffers from large performance gaps and strong forgetting, reflected by low accuracy and negative backward transfer. OPCM [70] improves over TA via orthogonalized merging but still shows notable degradation.
Figure 1: After 8-task continual merging: accuracy on first four tasks and overall BWT.
To overcome these limitations, we propose a novel continual merging paradigm—Test-Time Continual Model Merging (TTCMM)—which explicitly introduces the concept of test-time adaptation (TTA) [75, 63] into model merging. Unlike prior TTA-based multi-task merging methods [87, 68, 88], which assume simultaneous availability of models and test data from all tasks, TTCMM utilizes only a small set of unlabeled samples from the current task, making it uniquely suited for realistic continual scenarios where revisiting historical data is often infeasible.
In this paper, we propose Mingle (MIxture of Null-Space Gated Low-Rank Experts), a method designed to continually merge independently fine-tuned models at test-time while preserving prior knowledge. Mingle employs a mixture-of-experts architecture [27, 45] composed of lightweight LoRA-based [21] experts, enabling efficient and flexible test-time adaptation. To robustly prevent interference from previously learned tasks, we introduce a novel Null-Space Constrained Gating mechanism, restricting gating updates to task-orthogonal subspaces. Additionally, we propose an Adaptive Relaxation Strategy to dynamically modulate constraint strength based on test-time interference feedback during adaptation.
Extensive experiments on standard continual learning benchmarks show that Mingle consistently outperforms previous state-of-the-art approaches by 7–9% on average, achieving robust generalization and strong resistance to catastrophic forgetting across diverse continual learning scenarios. Remarkably, these improvements are achieved entirely without any access to original training data, demonstrating the effectiveness of our TTCMM paradigm and the power of test-time adaptation in continual learning.
Our contributions are summarized as follows:
  • We formalize test-time continual model merging (TTCMM), a novel task that leverages unlabeled test samples to merge independently fine-tuned models.
  • We propose Mingle, a TTCMM framework with Adaptive Null-Space Constrained Gating to effectively balance stability and plasticity.
  • Extensive experiments show that Mingle achieves state-of-the-art performance, consistently outperforming prior methods in accuracy, robustness and resistance to forgetting.
2 Related Work
Continual Learning. Continual learning (CL) seeks to mitigate catastrophic forgetting [43], where learning new tasks overwrites prior knowledge. Regularization-based methods constrain updates with importance weights [31, 92, 2, 30, 82], while distillation aligns outputs to preserve knowledge [20, 12, 60, 52]. Replay methods store exemplars or generate surrogates with prompts, prototypes, or generators [55, 39, 79, 61, 53], and dynamic architectures expand capacity via growth or ensembling [35, 98, 42]. Recent work leverages lightweight adapters or prompts in pre-trained models for efficient transfer [89, 23, 80]. Model merging offers an alternative route. Some methods remain close to conventional CL by sequentially fine-tuning and merging models to reduce forgetting [41, 42, 16], typically requiring training data. In contrast, continual model merging [28, 38, 5, 51, 70] merges independently fine-tuned models without revisiting training data, enabling greater scalability and privacy.
Model Merging. Early work merged models via direct parameter averaging [72, 58], later refined by linear mode connectivity [13, 1]. Wortsman et al. [81] showed that weight averaging can also enhance robustness and out-of-distribution generalization. Task Arithmetic (TA) [24] views models as task vectors to be summed, but relies on weight disentanglement [49], often violated under standard fine-tuning, motivating structured training [28, 65]. Beyond averaging, interference-aware methods reweight or sparsify parameters [86, 90], or fuse models via distillation and clustering [88, 76]. LoRA-based tuning [21] introduces additional entanglement challenges, spurring gradient-free or retrieval-based strategies [22, 94, 95]. More recently, dynamic merging [68, 40] adapts parameters conditioned on inputs, achieving higher flexibility and performance, but remains limited to multi-task fusion and unexplored in continual settings.
Figure 2: Comparison of three continual learning paradigms. (a) Conventional Continual Learning trains models sequentially with data arriving in stages, without access to previous task data. (b) Continual Model Merging continually fuses independently trained models, without access to any training data. (c) Test-Time Continual Model Merging improves merging by leveraging a few unlabeled test samples from the current task.
Test-Time Adaptation. TTA adapts models at inference to mitigate distribution shift. Early approaches used self-supervised objectives [66], entropy minimization [75], or regularized updates [63]. Online TTA adapts continuously [25], while batch-wise variants ignore temporal structure [17]. To enhance stability, later work introduced confidence filtering [48], EMA [11], partial updates [91], test-time augmentation [93], and adaptive BatchNorm [56]. For vision–language models, TTA often employs prompts or adapters [59, 15, 36]. We draw on TTA to guide merging, aligning fused models with evolving test distributions.
Relation to Prior Work. Most related to our work are MoE-Adapter [89] and WEMOE [68], both built on MoE architectures [27, 29]. MoE-Adapter follows conventional continual learning, embedding expert modules that are jointly trained across tasks. In contrast, we adopt a model-merging paradigm, where experts are extracted from independently fine-tuned models and inserted without further training. WEMOE incorporates test-time adaptation but targets multitask learning, assuming simultaneous access to all models and data. By contrast, Mingle is tailored for the more challenging continual merging setting.
3 Mingle: Mixture of Null-Space Gated Low-Rank Experts
3.1 Preliminaries
Problem Setting.
We study continual learning in a model merging setting, where a sequence of task-specific models $\{\theta_{1},\dots,\theta_{T}\}$ are independently fine-tuned from a shared pre-trained model $\theta_{0}$, each using a dataset $\mathcal{D}_{i}=\{(x_{j}^{(i)},y_{j}^{(i)})\}$ with label space $\mathcal{C}_{i}\subset\mathcal{Y}$. The goal is to construct a unified model $\theta^{\text{merged}}_{T}$ that generalizes across the combined label space $\mathcal{C}_{1:T}=\bigcup_{i=1}^{T}\mathcal{C}_{i}$.
Unlike conventional continual learning, we assume no access to training data during merging. All adaptation happens directly in parameter space. This paradigm is relevant in scenarios where only final fine-tuned models are retained, while original training data is discarded due to privacy, storage, or accessibility constraints.
To contextualize this, we compare with two related paradigms in Fig. 2:
  • Conventional Continual Learning. A single model $\theta$ is sequentially updated on $\mathcal{D}_{1},\dots,\mathcal{D}_{T}$, discarding previous data. It requires direct training data access and extensive retraining.
  • Continual Model Merging. A sequence of models are merged incrementally in parameter space without access to training data and earlier models: $\theta^{\text{merged}}_{t}=\text{Merge}(\theta^{\text{merged}}_{t-1},\theta_{t})$.
  • Test-Time Continual Model Merging. An extension of the above where a small unlabeled subset $\mathcal{D}^{\text{seed}}_{t}\subset\mathcal{D}^{\text{test}}_{t}$ (e.g., 5 samples per class) is available at each stage to provide lightweight task-specific guidance. We refer to $\mathcal{D}^{\text{seed}}_{t}$ as the seed samples of task $t$.
Existing Continual Merging Strategies. Let $\theta_{0}$ denote the parameters of a pre-trained model. The corresponding task vector is defined as $\Delta\theta_{t}=\theta_{t}-\theta_{0}$.
  • Continual Task Arithmetic (C. TA). A simple additive merge [24]: $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda\Delta\theta_{t}$, where $\lambda$ is a scalar. While training-free, it is sensitive to $\lambda$ and prone to task interference.
  • Orthogonal Projection-based Continual Merging (OPCM). Tang et al. [70] propose projecting each $\Delta\theta_{t}$ onto the orthogonal complement of previous directions: $\theta^{\text{merged}}_{t}=\theta_{0}+\frac{1}{\lambda_{t}}\left[\lambda_{t-1}\Delta\theta^{\text{merged}}_{t-1}+\mathcal{P}^{(t-1)}(\Delta\theta_{t})\right]$, where $\mathcal{P}^{(t-1)}$ retains components orthogonal to previous updates. This reduces interference but ignores adaptation to task distributions.
To address these issues, we present Mingle, which leverages $\mathcal{D}^{\text{seed}}_{t}$ to modulate the integration of $\theta_{t}$ at test-time, enhancing alignment to test distribution and mitigating task interference.
3.2 Motivation and Theoretical Analysis
Most existing continual model merging methods combine fine-tuned models via static averaging, where each expert is assigned fixed coefficients, thereby enforcing the same mixing rule across the whole input space. Consequently, it cannot specialize to regions where one expert is clearly superior. In contrast, a Mixture-of-Experts (MoE) equips every input with a data-dependent gate $g(x)=(g_{1}(x),\dots,g_{T}(x))$ that selects or re-weights experts on-the-fly. We give a formal comparison between static averaging and dynamic MoE under a noisy-routing scenario.  [^1] Symbols and proofs are deferred to App. A
Theorem 1 (Dynamic MoE versus Static Averaging).
Let $\{(D_{t},f_{t})\}_{t=1}^{T}$ be $T$ independent tasks with priors $P(t)$ and per-task risks $R_{t}(i)$. For any static mixture $h_{\mathrm{static}}(x)=\sum_{i=1}^{T}\alpha_{i}\,f_{i}(x)$ and any hard-routed MoE $h_{\mathrm{MoE}}(x)=f_{i^{\star}(x)}(x)$ with task-specific routing errors $\varepsilon_{t}$:
$R(h_{\mathrm{MoE}})\;=\;R_{\mathrm{ideal}}+\sum_{t=1}^{T}P(t)\,\varepsilon_{t}\bigl(R_{\text{wrong},t}-R_{t}(t)\bigr),$ (1)
where $R_{\mathrm{ideal}}=\sum_{t}P(t)R_{t}(t)$ and $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$. Moreover,
  1. 1.
    (Perfect routing) If $\varepsilon_{t}=0$ for all $t$, then $\inf_{g}R(h_{\mathrm{MoE}})\;<\;\inf_{\boldsymbol{\alpha}}R(h_{\mathrm{static}})$ whenever at least two tasks disagree on their best expert.
  2. 2.
    (Noisy routing)  If $\sum_{t}P(t)\varepsilon_{t}\bigl(R_{\text{wrong},t}-R_{t}(t)\bigr)<R^{*}_{\mathrm{static}}-R_{\mathrm{ideal}},$ where $R^{*}_{\text{static}}=\inf_{\boldsymbol{\alpha}}R(h_{\text{static}})$, then the MoE still attains lower risk than any static mixture.
The theory above motivates a design that (i) keeps experts specialized and (ii) prevents interference between tasks. Our Mingle framework achieves both goals by combining
  • Low-rank experts $f_{t}$ that capture task-specific variations with minimal parameters, and
  • Null-space constrained gating that projects gradient updates away from subspaces spanned by previously activated features, keeping $\varepsilon_{t}$ small without harming earlier experts.
3.3 Low-Rank Expert Mixture for Continual Model Merging
We adopt MoE framework for continual model merging, in which each task $i$ is equipped with a low-rank expert $f_{i}$ and an associated input-dependent gating function $g_{i}$. These components are injected into the linear layers of the backbone (e.g., CLIP visual encoder). The gate $g_{i}$ modulates expert activation based on the input features, allowing for fine-grained, localized task specialization.
Mixture of Low-Rank Expert. When a new task $t$ arrives, a dedicated expert $f_{t}$ and its gate $g_{t}$ are appended to the model. The output of a given $l$-th layer [^2] The layer index $l$ is omitted hereafter whenever it does not cause ambiguity. can be formulated as follows:
$\theta_{t}^{\text{merged},(l)}(X)=\theta_{t-1}^{\text{merged},(l)}(X)+g_{t}^{(l)}(X)\cdot f_{t}^{(l)}(X)=\theta_{0}^{(l)}(X)+\sum_{i=1}^{t}g_{i}^{(l)}(X)\cdot f_{i}^{(l)}(X).$ (2)
where only the gate $g_{t}$ is adaptable during testing, while all experts $\{f_{i}\}_{i=1}^{t}$ and old gates $\{g_{i}\}_{i=1}^{t-1}$ remain frozen to preserve prior knowledge. To construct expert $f_{t}$, we first project the task vector $\Delta\theta_{t}$ onto the orthogonal complement of previously learned directions, following OPCM [70]:
$\mathcal{P}^{(t-1)}(\Delta\theta_{t})=\sum_{\begin{subarray}{c}p,q=a,p\neq q\end{subarray}}^{m,n}\langle\Delta\theta_{t},u_{p}^{(t-1)}v_{q}^{(t-1)\top}\rangle_{F}u_{p}^{(t-1)}v_{q}^{(t-1)\top},$ (3)
where $u_{p}^{(t-1)}$ and $v_{q}^{(t-1)}$ are the $p$-th and $q$-th singular vectors from the singular value decomposition (SVD) of previous experts $\sum_{i=1}^{t-1}f_{i}(X)$, and $\alpha$ denotes the effective rank of previous experts. This projection removes previously learned directions to mitigate interference. We then apply a rank-$r$ truncated SVD for $\mathcal{P}^{(t-1)}(\Delta\theta_{t})$ to construct a low-rank expert [21]:
$f_{t}=BA=(\tilde{U}\tilde{\Sigma})\tilde{V}^{\top}$, (4)
where $\tilde{U}\in\mathbb{R}^{d_{1}\times r}$, $\tilde{\Sigma}\in\mathbb{R}^{r\times r}$, and $\tilde{V}\in\mathbb{R}^{d_{2}\times r}$, retaining the top $r$ singular components. The resulting expert captures the principal directions while significantly reducing parameter overhead.
Each gating function is implemented as a linear projection:
$g_{t}(X)=W_{t}^{(g)\top}X+b_{t}^{(g)},$ (5)
where $W_{t}^{(g)}\in\mathbb{R}^{d\times 1}$ and $b_{t}^{(g)}\in\mathbb{R}$ are learnable parameters. The gating function is adapted at test time using a small number of unlabeled test data.
Test-Time Adaptation. To encourage the merged model to retain task-specific behavior, we minimize the Kullback–Leibler divergence between its prediction and that of the corresponding individual fine-tuned model $\theta_{t}$. We define the adaptation objective as:
$L_{t}=\mathbb{E}_{x\sim\mathcal{D}^{\text{seed}}_{t}}\left[\mathrm{KL}\left(p(x;\;\theta^{\text{merged}}_{t})\,\|\,p(x;\;\theta_{t})\right)\right].$ (6)
where $p(x;\theta)$ denotes the predictive distribution
3.4 Adaptive Null-Space Constrained Gating for Interference Mitigation
When merging models continually, the primary challenge of gating is to integrate new experts without disturbing prior task predictions. Consider two experts $f_{1},f_{2}$ and their corresponding gates $g_{1},g_{2}$. When evaluating on the first task domain $X_{1}$, the interference from $g_{2}$ can be quantified as:
$\xi(g_{2})=\left\|g_{1}(X_{1})\cdot f_{1}(X_{1})+g_{2}(X_{1})\cdot f_{2}(X_{1})-f_{1}(X_{1})\right\|^{2}.$ (7)
This measures the deviation introduced by $g_{2}$ on the domain where $f_{1}$ originally dominates. A desirable gating function should suppress $g_{2}(X_{1})$, ensuring predictions on $X_{1}$ remain unaffected. However, as $X_{1}$ becomes inaccessible after adaptation, this error becomes unobservable and cannot be minimized directly, resulting in prediction drift and catastrophic forgetting.
Hard Null-Space Projection. After completing task $t$, we cache the $l$-th layer inputs in the seed buffer $\mathcal{D}_{t}^{\text{seed}}$ and estimate their covariance $\mathrm{Cov}_{t}^{(l)}\in\mathbb{R}^{d\times d}$. Applying truncated SVD yields the top-$k$ dominant subspaces $\tilde{U}_{t}^{(l)}\in\mathbb{R}^{d\times k}$. We then concatenate these with the subspaces from all previous tasks and orthonormalize: $U_{t}^{(l)}=\operatorname{orthonorm}\bigl[U_{t-1}^{(l)}|\tilde{U}_{t}^{(l)}\bigr]\in\mathbb{R}^{d\times tk}$. The hard projector is $P_{t}=I-U_{t}^{(l)}\,U_{t}^{(l)\top}\in\mathbb{R}^{d\times d}.$ To suppress interference from tasks $\leq t-1$, the gating update for task $t$ is $W_{t}^{(g,l)}\leftarrow W_{t}^{(g,l)}-\eta\,\nabla L_{t}^{(l)}P_{t-1}^{(l)}$. However, this projection may also discard gradient components that are informative for task $t$ whenever its feature support overlaps with $\operatorname{span}(U_{t-1})$.
Input: pre-trained model $\theta_{0}$ and fine-tuned models $\{\theta_{t}\}_{t=1}^{T}$; seed data $\{\mathcal{D}^{\text{seed}}_{t}\}_{t=1}^{T}$; hyper-parameters $k,\beta,\gamma$; learning rate $\eta$
Output: Merged model $\theta_{T}^{\text{merged}}$
Init: $\theta_{0}^{\text{merged}}\!\leftarrow\!\theta_{0}$;
$U_{0}^{(l)}\!\leftarrow\!\emptyset$
for task $t=1$ to $T$ do
 $\triangleright$ create low-rank experts (Eqs. 3 and 4)
 $f_{t}={\small\textsc{SVD}_{\textsc{trunc}.}}~\!\bigl(\,\underbrace{\mathcal{P}^{(t-1)}(\theta_{t}-\theta_{0})}_{\scriptsize\text{use }\theta_{1}-\theta_{0}\text{ when }t=1}\bigr)=B_{t}A_{t}$
 $\triangleright$ add expert & initialize gate (Eq. 5)
 $\theta_{t}^{\text{merged}}=\theta_{t-1}^{\text{merged}}+g_{t}\!\cdot\!f_{t},\;\{W_{t}^{(g)},b_{t}^{(g)},S^{(0)}\}\leftarrow 0$
 for $m=1$ to total iterations do
    $X\leftarrow\text{batch}(\mathcal{D}^{\text{seed}}_{t})$
    $\{\nabla_{W_{t}^{(g)}}L,\nabla_{b_{t}^{(g)}}L\}\leftarrow\nabla L_{t}(X,\theta_{t}^{\text{merged}},\theta_{t})$
    if $t>1$ then
       $\triangleright$ project gradient onto null-space (Eqs. 8-11)
       $S^{(m)}=\beta S^{(m-1)}+(1-\beta)r$
       $\nabla_{W_{t}^{(g)}}L_{t}\leftarrow\nabla_{W_{t}^{(g)}}L_{t}\,\widetilde{P}_{t-1}$
      end if
    $W_{t}^{(g)}\leftarrow W_{t}^{(g)}-\eta\,\nabla_{W_{t}^{(g)}}L_{t}$
    $b_{t}^{(g)}\leftarrow b_{t}^{(g)}-\eta\,\nabla_{b_{t}^{(g)}}L_{t}$
   end for
 $\triangleright$ update dominant subspaces
 $U_{t}\leftarrow\operatorname{orthonorm}\!\bigl[U_{t-1}\mid\tilde{U}_{t}]$
end for
Algorithm 1 Mingle Procedure.
Adaptive Null-Space Relaxation. To restore plasticity, we replace the all-one eigenvalues of $P_{t-1}$ with soft coefficients learned online.
(i) Interference statistics. For each column $u_{p}^{(l)}$ of $U_{t-1}^{(l)}$ we measure instantaneous alignment:
$r_{p}^{(l)}=\|(\nabla L_{t})^{\top}u_{p}^{(l)}\|_{2}\;/\;\|\nabla L_{t}\|_{2}.$ (8)
We maintain per-direction interference scores $S^{(m,l)}\in\mathbb{R}^{k}$ (initialized to 0) by applying an exponential moving average at each iteration $m$:
$S^{(m,l)}=\beta\,S^{(m-1,l)}+(1-\beta)\,r^{(l)},$ (9)
which suppresses stochastic gradient noise while preserving the dominant interference directions.
(ii) Adaptive shrinkage. Each direction is attenuated by $\lambda^{(m,l)}=\exp(-\gamma\,S^{(m,l)})$ ($\gamma>0,\ \lambda^{(m,l)}\in(0,1]$). Let $\Lambda^{(m,l)}_{t-1}=\mathrm{diag}(\lambda_{1}^{(m,l)},\dots,\lambda_{(t-1)\cdot k}^{(m,l)})$. The relaxed projector becomes:
$\widetilde{P}_{t-1}^{(m,l)}=U_{t-1}^{(m,l)}\,\Lambda^{(m,l)}_{t-1}\,U_{t-1}^{(m,l)\top},$ (10)
interpolating smoothly between no protection ($\Lambda=0$) and the hard null projector ($\Lambda=I$).
(iii) Update rule. We finally update:
$W_{t}^{(g,l)}\;\leftarrow\;W_{t}^{(g,l)}-\eta\,\nabla L_{t}^{(l)}\,\widetilde{P}_{t-1}^{(l)}.$ (11)
Relaxing the projector inevitably allows more residual interference than the hard null‐space variant, yet empirically the increase is minor and is offset by markedly higher plasticity (Tab. 5), indicating a favorable stability–plasticity balance. The overall procedure is outlined in Algo. 1.
4 Experiments
We describe the experimental setup in Sec. 4.1, followed by the main results in Sec. 4.2 and further analysis and ablations in Sec. 4.3. Due to page limitations, detailed results are provided in the Appendix.
4.1 Experimental Setup
Datasets and Models.  Following [70], we evaluate on image-classification tasks with CLIP-ViT backbones [54]. We consider 8, 14, and 20-task groups using ViT-B/32, ViT-B/16, and ViT-L/14 models, each fine-tuned on up to 20 downstream tasks, with checkpoints from FusionBench [67]. To assess order sensitivity, we repeat experiments over 10 random seeds (42–51). For comparison with conventional CL, we use the Multi-domain Task-Incremental Learning (MTIL) benchmark [97] with eleven vision tasks. Beyond vision, we evaluate on eight GLUE language tasks [74] with a Flan-T5-base backbone [6].
Implementation Details.  We insert low-rank experts into the CLIP vision encoder. Two variants are used: a full setup modifying all attention and MLP layers, and a lightweight one on attn.qkv and mlp.fc1. All experiments share a single set of global hyper-parameters across models and task orders. Each expert has rank $r=64$; the null-space constraint uses $k=3$, $\gamma=1$, and $\beta=0.99$. Adaptation runs for 50 iterations with Adam (lr $1\text{e-4}$, batch size 16). For vision tasks we use 5 unlabeled samples per class, and for NLP tasks 100 in total, all without access to prior-task data.
Evaluation Metrics. We evaluate using average accuracy (ACC) and backward transfer (BWT) [37]. ACC is the mean accuracy of the final merged model across all tasks: $\text{ACC}=\tfrac{1}{T}\sum_{i=1}^{T}a_{i}(\theta_{t}^{\text{merged}})$, where $a_{i}(\cdot)$ is accuracy on task $i$. BWT measures forgetting by comparing performance on earlier tasks before vs. after the final merge: $\text{BWT}=\tfrac{1}{T-1}\sum_{i=1}^{T-1}\big[a_{i}(\theta_{T}^{\text{merged}})-a_{i}(\theta_{i}^{\text{merged}})\big]$.
[…]
Method Assump. ViT-B/32 ViT-B/16 ViT-L/14
DM / DA 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
Pre-Trained –  /  – 48.1 56.9 55.6 55.4 62.0 59.8 64.9 69.1 65.6
Fine-Tuned –  /  – 90.4 89.3 89.8 92.4 91.3 91.6 94.3 93.4 93.5
C. Fine-Tuned –  /  – 79.8 67.4 62.6 82.9 72.2 68.2 90.0 70.9 77.7
ACC (%) $\uparrow$ Average (SWA) [26]  /  66.3 $\pm$0.0 65.4 $\pm$0.0 61.1 $\pm$0.0 72.3 $\pm$0.0 69.7 $\pm$0.0 64.8 $\pm$0.0 80.0 $\pm$0.0 77.5 $\pm$0.0 71.1 $\pm$0.0
C. Task Arithmetic [24]  /  67.5 $\pm$0.0 66.5 $\pm$0.0 60.0 $\pm$0.0 77.1 $\pm$0.0 70.9 $\pm$0.6 64.2 $\pm$0.0 82.1 $\pm$0.0 77.9 $\pm$0.0 70.3 $\pm$0.0
C. Ties-Merging [86]  /  49.0 $\pm$10.2 66.2 $\pm$0.6 59.9 $\pm$0.7 66.8 $\pm$3.7 70.5 $\pm$0.8 63.0 $\pm$1.6 64.3 $\pm$7.0 78.0 $\pm$0.6 68.3 $\pm$0.9
MagMax-Ind [41]  /  70.7 $\pm$0.0 67.0 $\pm$0.0 61.2 $\pm$0.0 76.7 $\pm$1.8 67.0 $\pm$0.0 62.5 $\pm$0.0 83.4 $\pm$0.0 71.2 $\pm$0.0 71.2 $\pm$0.0
Consensus TA [77]  /  67.1 $\pm$0.4 64.1 $\pm$0.8 45.8 $\pm$1.5 72.8 $\pm$0.5 69.0 $\pm$0.0 49.9 $\pm$1.9 80.4 $\pm$0.5 75.0 $\pm$1.0 51.3 $\pm$2.4
OPCM [70]  /  75.5 $\pm$0.5 71.9 $\pm$0.3 65.7 $\pm$0.2 81.8 $\pm$0.3 77.1 $\pm$0.5 70.3 $\pm$0.2 87.0 $\pm$0.4 83.5 $\pm$0.2 76.0 $\pm$0.2
C. LW AdaMerging [87]  /  53.4 $\pm$ 3.2 59.8 $\pm$ 1.6 59.7 $\pm$ 7.4 59.9 $\pm$2.3 64.3 $\pm$ 1.2 61.5 $\pm$ 1.1 68.8 $\pm$ 2.9 73.1 $\pm$ 5.7 66.9 $\pm$ 1.1
C. LoRA-WEMOE [68]  /  68.8 $\pm$ 7.8 63.8 $\pm$ 3.4 49.6 $\pm$ 15.4 72.6 $\pm$ 3.7 67.9 $\pm$ 2.9 55.0 $\pm$ 7.0 75.6 $\pm$ 7.8 74.0 $\pm$ 5.0 56.9 $\pm$ 19.8
Mingle (Ours)  /  85.8 $\pm$0.8 81.6 $\pm$1.4 77.1 $\pm$2.0 88.3 $\pm$0.6 84.9 $\pm$0.8 81.9 $\pm$0.9 91.8 $\pm$0.2 88.8 $\pm$0.7 85.5 $\pm$1.3
Mingle (Ours)  /  85.0 $\pm$0.5 81.7 $\pm$1.0 77.1 $\pm$1.3 87.0 $\pm$0.6 84.7 $\pm$1.0 81.6 $\pm$1.3 91.4 $\pm$0.3 89.2 $\pm$0.1 83.6 $\pm$0.6
BWT (%) $\uparrow$ Average (SWA) [26]  /  -11.5 $\pm$2.2 -8.0 $\pm$1.3 -7.1 $\pm$2.1 -9.7 $\pm$1.5 -7.1 $\pm$1.4 -7.3 $\pm$1.7 -7.3 $\pm$1.4 -5.8 $\pm$1.0 -6.4 $\pm$1.5
C. Task Arithmetic [24]  /  -9.6 $\pm$1.5 -1.3 $\pm$1.6 -3.4 $\pm$1.0 -4.2 $\pm$1.0 -1.3 $\pm$0.4 -3.6 $\pm$0.4 -7.1 $\pm$0.8 -1.8 $\pm$0.3 -3.3 $\pm$0.3
C. Ties-Merging [86]  /  -15.3 $\pm$8.0 1.9 $\pm$0.6 -1.5 $\pm$0.7 -5.5 $\pm$0.4 1.4 $\pm$0.7 -1.5 $\pm$1.2 -13.0 $\pm$5.7 -1.1 $\pm$0.4 -2.9 $\pm$1.0
MagMax-Ind [41]  /  -8.3 $\pm$1.3 -7.4 $\pm$1.4 -7.2 $\pm$1.6 -6.1 $\pm$1.3 -7.4 $\pm$2.0 -8.0 $\pm$2.2 -5.0 $\pm$0.8 -6.0 $\pm$2.1 -6.5 $\pm$2.1
Consensus TA [77]  /  3.8 $\pm$0.9 -1.3 $\pm$0.9 -11.8 $\pm$1.9 3.5 $\pm$0.6 -1.1 $\pm$0.8 -11.6 $\pm$1.3 2.4 $\pm$0.6 -2.5 $\pm$0.8 -16.5 $\pm$1.5
OPCM [70]  /  -6.3 $\pm$1.1 -6.0 $\pm$1.0 -7.8 $\pm$1.5 -4.8 $\pm$0.7 -5.1 $\pm$1.4 -6.3 $\pm$2.2 -2.6 $\pm$1.0 -4.3 $\pm$0.7 -6.5 $\pm$1.8
C. LW AdaMerging [87]  /  -32.5 $\pm$3.6 -24.1 $\pm$1.7 -22.7 $\pm$4.3 -27.8 $\pm$2.7 -22.1 $\pm$1.4 -21.4 $\pm$1.2 -24.3 $\pm$3.3 -19.6 $\pm$1.7 -21.7 $\pm$1.1
C. LoRA-WEMOE [68]  /  -20.4 $\pm$9.0 -20.2 $\pm$3.9 -24.5 $\pm$10.0 -18.0 $\pm$6.2 -18.8 $\pm$3.4 -25.8 $\pm$7.9 -17.8 $\pm$5.9 -16.8 $\pm$5.3 -27.9 $\pm$17.2
Mingle (Ours)  /  -0.6 $\pm$0.4 -1.1 $\pm$0.3 -2.2 $\pm$0.8 -0.4 $\pm$0.1 -0.9 $\pm$0.1 -1.9 $\pm$0.4 -0.6 $\pm$0.1 -1.0 $\pm$0.3 -2.6 $\pm$0.9
Mingle (Ours)  /  -0.1 $\pm$0.1 -0.4 $\pm$0.1 -1.3 $\pm$0.6 -0.1 $\pm$0.1 -0.3 $\pm$0.1 -1.0 $\pm$0.4 -0.2 $\pm$0.0 -0.4 $\pm$0.2 -1.5 $\pm$0.6
Table 2: Results of continual merging Flan-T5-base models on 8 tasks, ordered alphabetically.
Method DM / DA CoLA MNLI MRPC QNLI QQP RTE SST2 STSB ACC $\uparrow$ BWT $\uparrow$
Pre-trained –  /  – 69.1 56.5 76.2 88.4 82.1 80.1 91.2 62.2 75.7 -
Individual –  /  – 75.0 83.4 87.5 91.5 85.4 85.9 93.6 88.7 86.4 -
Task Arithmetic  /  69.1 58.1 77.9 88.9 83.1 79.1 90.7 74.0 77.6 -4.6
Ties-Merging  /  39.3 70.0 82.4 88.8 81.8 75.8 89.7 76.8 75.6 -6.1
OPCM  /  69.9 72.9 78.7 90.3 83.8 83.0 92.2 73.7 80.6 -2.5
LW AdaMerging  /  69.1 58.1 77.9 88.9 83.1 79.1 90.7 74.2 77.6 -4.7
LoRA-WEMOE  /  71.5 80.6 78.2 90.3 82.7 80.5 91.3 76.2 81.4 0.1
Mingle (Ours)  /  75.0 78.2 86.0 90.9 84.2 80.5 92.5 78.8 83.3 0.1
4.2 Main Results
Overall Performance  As shown in Tab. 1, Mingle substantially outperforms previous continual merging methods on all CLIP backbones and task counts. It achieves the highest accuracy with backward forgetting kept near zero, demonstrating both strong forward learning and long-term stability. The lightweight variant performs on par with the full version, further underscoring the robustness of our approach. On NLP benchmarks (Tab. 2), Mingle likewise attains the best overall accuracy and non-negative BWT, improving on multiple GLUE tasks while maintaining balanced performance across the suite. Together, these results across vision and language confirm that Mingle consistently delivers state-of-the-art accuracy while nearly eliminating forgetting under diverse continual scenarios.
Table 3: Comparison of last accuracy (%) with conventional CL approaches on MTIL benchmark.
Method
Aircraft
Caltech101
CIFAR100
DTD
EuroSAT
Flowers
Food101
MNIST
Pets
Cars
SUN397
 Avg.
Conventional CL (Sequential fine‐tuned)
WiSE-FT [81] 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
ZSCL [97] 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
MoE-Adapter [89] 49.8 92.2 86.1 78.1 95.7 94.3 89.5 98.1 89.9 81.6 80.0 85.0
DIKI [71] 45.2 95.7 86.3 72.9 98.0 97.0 89.2 99.4 94.2 81.6 76.6 85.1
AwoForget [96] 42.4 92.7 83.2 73.2 97.0 91.8 92.2 99.1 93.9 87.4 82.6 85.0
Dual-RAIL [85] 52.5 96.8 83.3 80.1 96.4 99.0 89.9 98.8 93.5 85.5 79.2 86.8
MagMax [41] 40.2 96.1 81.1 72.0 97.8 76.3 88.4 99.2 93.0 70.5 68.9 80.3
Mingle-Seq 58.7 97.5 87.2 79.7 97.3 87.2 90.1 99.6 93.0 80.4 73.3 85.8
Continual Merging (Independent fine‐tuned)
Average (SWA) [26] 26.5 92.3 74.3 48.4 73.7 74.0 87.1 84.0 91.2 67.5 68.5 71.6
C. TA [24] 26.6 92.5 74.5 48.7 74.3 74.4 87.0 85.5 91.2 67.7 68.6 71.9
C. Ties [86] 30.5 94.0 74.8 49.8 71.7 73.8 87.3 81.5 90.6 67.0 67.9 71.7
MagMax-Ind [41] 29.9 93.7 78.4 46.1 58.3 68.1 86.8 82.8 91.4 62.7 69.3 69.8
OPCM [70] 35.7 95.9 77.0 54.6 90.3 76.4 87.1 96.3 93.3 70.1 70.5 77.0
Mingle 54.2 97.3 79.7 72.3 96.0 86.7 88.7 99.3 93.9 73.1 71.6 83.0
Table 4: Robustness results of ViT-B/32 continually merged across 4 tasks.
Method Clean Motion Impulse Gaussian Pixelate Spatter Contrast JPEG Avg.
ACC (%) $\uparrow$ C. LW AdaMerging [87] 56.0 $\pm$5.3 47.5 $\pm$4.4 43.1 $\pm$2.3 43.3 $\pm$3.4 18.1 $\pm$4.7 46.6 $\pm$3.0 48.9 $\pm$4.8 49.1 $\pm$4.0 44.9
C. WEMOE [68] 3.4 $\pm$0.8 3.1 $\pm$0.4 4.3 $\pm$1.4 3.4 $\pm$1.4 3.0 $\pm$1.6 4.0 $\pm$0.9 3.3 $\pm$0.7 4.0 $\pm$1.2 3.6
C. LoRA-WEMOE [68] 78.7 $\pm$4.5 71.0 $\pm$4.9 55.0 $\pm$3.8 59.4 $\pm$3.8 24.9 $\pm$24.9 60.5 $\pm$3.8 68.5 $\pm$4.8 69.7 $\pm$4.4 61.0
C. Task Arithmetic [24] 77.5 $\pm$0.0 66.0 $\pm$0.0 58.9 $\pm$0.0 59.6 $\pm$0.0 29.7 $\pm$0.0 63.5 $\pm$0.0 66.0 $\pm$0.0 67.8 $\pm$0.0 61.1
MagMax-Ind [41] 79.1 $\pm$0.0 69.0 $\pm$0.0 60.6 $\pm$0.0 61.5 $\pm$0.0 33.0 $\pm$0.0 66.4 $\pm$0.0 68.6 $\pm$0.0 69.9 $\pm$0.0 63.5
OPCM [70] 83.6 $\pm$0.5 72.5 $\pm$0.6 64.7 $\pm$1.2 65.2 $\pm$1.2 35.2 $\pm$0.6 70.5 $\pm$0.5 72.5 $\pm$0.6 74.4 $\pm$0.3 67.3
Mingle (Ours) 89.9 $\pm$0.4 82.8 $\pm$0.8 67.5 $\pm$2.0 70.7 $\pm$1.2 37.9 $\pm$0.4 77.0 $\pm$0.7 80.1 $\pm$0.8 82.9 $\pm$0.9 73.2
BWT (%) $\uparrow$ C. LW AdaMerging [87] -38.0 $\pm$7.1 -37.3 $\pm$5.9 -22.2 $\pm$3.0 -25.2 $\pm$4.5 -20.8 $\pm$6.3 -28.6 $\pm$4.0 -34.7 $\pm$6.5 -36.1 $\pm$5.3 -29.5
C. WEMOE [68] -30.7 $\pm$3.1 -28.7 $\pm$3.8 -22.1 $\pm$11.6 -25.5 $\pm$9.8 -8.0 $\pm$4.5 -23.4 $\pm$11.0 -27.6 $\pm$5.6 -28.6 $\pm$5.2 -24.3
C. LoRA-WEMOE [68] -13.6 $\pm$6.9 -14.6 $\pm$8.2 -11.3 $\pm$4.7 -10.2 $\pm$3.7 -15.6 $\pm$8.8 -10.8 $\pm$3.9 -11.2 $\pm$8.8 -16.7 $\pm$6.6 -13.0
C. Task Arithmetic [24] -4.8 $\pm$0.9 -6.1 $\pm$1.2 -1.6 $\pm$3.0 -1.6 $\pm$1.7 -2.7 $\pm$1.5 -3.1 $\pm$2.5 -6.1 $\pm$1.2 -5.1 $\pm$0.8 -3.9
MagMax-Ind [41] -7.7 $\pm$0.8 -8.1 $\pm$1.5 -6.1 $\pm$4.9 -5.1 $\pm$3.7 -3.5 $\pm$3.0 -7.3 $\pm$2.9 -8.4 $\pm$1.6 -8.2 $\pm$1.2 -6.8
OPCM [70] -4.3 $\pm$1.8 -4.5 $\pm$2.8 -6.4 $\pm$7.1 -6.1 $\pm$4.3 -2.9 $\pm$0.9 -6.3 $\pm$2.9 -4.5 $\pm$2.8 -5.7 $\pm$1.5 -5.1
Mingle (Ours) -0.2 $\pm$0.2 -0.1 $\pm$0.4 0.7 $\pm$1.0 0.6 $\pm$1.1 -0.2 $\pm$1.1 0.2 $\pm$0.7 0.0 $\pm$0.5 0.5 $\pm$0.5 -0.2
Comparison with Conventional CL. Tab. 3 evaluates two CL paradigms on the MTIL benchmark: conventional CL, where each task model is fine-tuned from its immediate predecessor; and continual merging, which fine-tunes each task model independently before fusion, eliminating inter-model dependencies and enabling flexible task ordering and model reuse. Within the merging family, Mingle sets a new state‑of‑the‑art, and when integrated into a sequential fine‑tuning pipeline, it matches the performance of SOTA CL methods. This demonstrates both its strength as a fusion strategy and its versatility across different training regimes.
Robustness to Test-Time Distribution Shifts. Following prior work [87, 68], we evaluate Mingle on seven corruptions (motion blur, impulse noise, Gaussian noise, pixelate, spatter, contrast, JPEG) and report results in Tab. 4. It preserves high accuracy and near‑zero or even positive BWT, outperforming all baselines, whereas direct application of SOTA TTA‑based merging (WEMOE, AdaMerging) in a continual setting fails without tailored designs to continual setup.
4.3 Ablation Results and Analysis
Ablation Study. We explore the contribution of each component in Tab. 5. Row 1 shows a fixed‑weight merging of low‑rank experts as our baseline. In Row 2, adding TTA boosts ACC substantially but at the cost of worsening BWT with more tasks. Row 3 demonstrates that freezing earlier gates curbs forgetting while retaining ACC gains. Row 4 then applies null‑space constraints, yielding further BWT improvements. Finally, Row 5 presents the full method with adaptive relaxation, which best harmonizes accuracy and long‑term stability.
Table 5: Ablation study of Mingle with CLIP ViT-B/16 over 8, 14, and 20 tasks.
Test-Time Frozen Null-Space Adaptive ACC(%) BWT(%)
Adaptation Old Gate Constrained Gate Relaxation 8 tasks 14 tasks 20 tasks 8 tasks 14 tasks 20 tasks
78.7 $\pm$0.1 76.4 $\pm$1.0 70.6 $\pm$0.4 -0.5 $\pm$0.1 -1.0 $\pm$0.3 -1.3 $\pm$0.3
86.4 $\pm$5.3 81.7 $\pm$2.3 76.7 $\pm$1.3 -6.0 $\pm$5.2 -7.7 $\pm$-3.4 -12.8 $\pm$1.3
87.4 $\pm$0.4 81.3 $\pm$0.8 76.2 $\pm$1.3 -2.3 $\pm$0.5 -4.3 $\pm$0.8 -6.8 $\pm$0.9
86.0 $\pm$1.5 83.5 $\pm$0.9 78.3 $\pm$1.7 -0.1 $\pm$0.1 -0.1 $\pm$0.1 -0.2 $\pm$0.1
88.3 $\pm$0.6 84.9 $\pm$0.8 81.9 $\pm$0.9 -0.4 $\pm$0.1 -0.9 $\pm$0.1 -1.9 $\pm$0.4
Table 6: Ablation on number of adaptation steps for ViT-B/32 across 8, 14, and 20 tasks.
Steps ACC (8-task) BWT (8-task) ACC (14-task) BWT (14-task) ACC (20-task) BWT (20-task)
[…]
Table 7: Efficiency and sample analysis. (a) Expert insertion layers and rank sweep over 8 tasks on CLIP-ViT-B/32. (b) Wall-clock adaptation time across tasks on CLIP-ViT-B/32. (c) Accuracy (%) of CLIP-ViT-B/16 under varying numbers of test samples.
(a) Expert insertion layers and rank sweep (8 tasks)
Configuration TTA Time Train. Param Full Param ACC(%)
attn.qkv_proj ($r=64$) 61 s 27.7 k 116.0 M 69.9
attn.out_proj ($r=64$) 47 s 9.3 k 97.0 M 53.9
mlp.fc1 ($r=64$) 48 s 9.0 k 111.1 M 82.6
mlp.fc2 ($r=64$) 48 s 36.8 k 111.3 M 70.1
attn & mlp ($r=64$) 78 s 83.0 k 173.1 M 85.8
qkv & fc1 ($r=32$) 65 s 36.9 k 113.7 M 84.5
qkv & fc1 ($r=128$) 65 s 36.9 k 191.6 M 85.1
qkv & fc1 ($r=768$) 70 s 36.9 k 710.3 M 83.7
qkv & fc1 ($r=64$) 65 s 36.9 k 139.7 M 85.0
(b) Wall-clock adaptation time across tasks
#Tasks Adaptation steps Total Time (s) Avg./task (s)
8 50 78 9.8
14 50 138 9.9
20 50 211 10.6
(c) Results with varying samples/class
# Samples/class 8-task 14-task 20-task
[…]
Ablation on TTA optimization steps. Tab. 6 reports accuracy and backward transfer of CLIP ViT-B/32 under different adaptation steps across 8, 14, and 20 tasks. Accuracy improves steadily with longer schedules, rising from 60–63% at 5 steps to 77–86% at 50 steps. Forgetting remains negligible, with BWT close to zero in all cases and only a minor drop of about 2% in the 20-task setting at 50 steps. Notably, as few as 20 steps are sufficient to surpass all baselines in Tab. 1, while 50 steps yield the best accuracy with only a modest increase in cost. These results confirm that Mingle is both effective under tight compute budgets and scalable with additional adaptation.
Computation and Parameter Efficiency. Tab. 7 summarizes the efficiency analysis. In part (a), inserting experts into both attn and mlp layers yields the highest accuracy, 85.8%, but also the longest adaptation time of 78s and 83k trainable parameters. A lighter hybrid scheme qkv & fc1 with rank 64 reaches 85.0% accuracy with 36.9k parameters and 65s, offering a better trade-off. The rank sweep shows that raising the rank from 32 to 64 improves accuracy from 84.5% to 85.0%, while larger ranks bring little or even negative gain, e.g., rank 768 drops to 83.7%. Part (b) reports wall-clock adaptation time as tasks increase: with 20 tasks the total is 211s, averaging about 10s per task. After adaptation the router remains fixed and inference is purely feedforward without TTA, enabling low-latency deployment across all tasks. Overall, Mingle achieves a strong balance of accuracy, parameter efficiency, and scalability under diverse resource budgets.
Number of Seed Samples. The number of seed samples per class is crucial for TTA reliability and efficiency. As shown in Tab. 7 (c), using no samples reduces to the Static baseline, where LoRA modules are merged with fixed coefficients (0.3) without adaptation, yielding 70–79% accuracy. Introducing a single sample lifts accuracy to 81–88%, and adding more samples offers only minor gains. Variance across task orders decreases with more samples, making five samples per class a balanced trade-off between performance and efficiency.
Figure 3: Gate activations across eight tasks under varying $\gamma$. Each subplot shows one gate; curves and shaded areas indicate mean and std across layers. Gray bars mark the gate’s training task. Lower $\gamma$ leads to stronger suppression on prior tasks.
(a) Vary $\gamma$, $\beta=0.99$, $k=3$
(b) Vary $\beta$, $\gamma=1$, $k=3$
(c) Vary $k$, $\gamma=0$
Figure 4: Sensitivity analysis of the null-space constrained gating w.r.t. hyper-parameters $\beta$, $\gamma$, and $k$.
Visualization on Gate Activations. Fig. 3 shows that the null-space constraint suppresses gate responses on previously seen tasks, reducing forgetting, with smaller $\gamma$ giving stronger attenuation. We also observe that gate activations remain below 1.0 across tasks, even in the w/o Null-Space variant (blue curve), showing that under-activation is not solely due to the constraint. Instead, it reflects the complementary nature of experts: multiple LoRA modules capture overlapping but distinct subspaces, and softly combining them often yields better performance, especially in settings like TTCMM where task boundaries are fuzzy.
Hyper-parameter of Gate. We study the effect of three key hyper-parameters: $\gamma$, $\beta$, and $k$. $\gamma$ controls the strength of null-space suppression; smaller values lead to stronger attenuation of activations on prior tasks, reducing forgetting (Fig. 4(a)). As shown in Fig. 4(b), $\beta$ regulates the smoothness of the EMA used to accumulate interference signals. A moderate setting ($\beta=0.99$) balances responsiveness and stability. Smaller $\beta$ amplifies noise sensitivity, while larger $\beta$ slows detection of interference. $k$ determines the number of principal directions retained per task. The mitigation of forgetting saturates at $k=3$ (Fig. 4(c)), indicating that a small number of task-specific directions suffices.
5 Conclusions
In this work, we introduced the task of test-time continual model merging (TTCMM) and proposed Mingle, a novel framework for TTCMM that integrates a mixture-of-experts architecture with adaptive null-space constrained gating. Extensive empirical evaluations show that Mingle substantially improves generalization and mitigates catastrophic forgetting, consistently outperforming prior state-of-the-art approaches. These results establish TTCMM as a principled paradigm for addressing both task interference and distribution shift, and highlight the practical potential of Mingle for scalable and efficient continual learning in real-world applications.
Acknowledgments
This work was supported in part by National Science and Technology Major Project (2021ZD0112001), National Natural Science Foundation of China (No.62271119, 08120002, 62071086, U23A20286), the Key Research and Development Project of Hainan Province (Grant No. ZDYF2024(LALH)003), the Fundamental Research Funds for the Central University of China (DUT No. 82232031), the Natural Science Foundation of Sichuan Province under Grant 2025ZNSFSC0475.
We thank all reviewers for taking the time to review our paper and give valuable suggestions.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the main contributions, which are consistently supported by both theoretical insights and empirical evaluations presented throughout the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the proposed approach are explicitly discussed in Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theoretical claims are accompanied by clear assumptions and complete proofs in Appendix, with theorem formally stated and rigorously derived.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Detailed descriptions of datasets, training protocols, and evaluation metrics are provided in Section 4.1 and Appendix, enabling faithful reproduction of the main results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification:The code and data are not provided at submission time, but the authors state they will release them upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper includes detailed descriptions of datasets, data splits, training procedures, hyperparameters, and evaluation protocols in both the main text and appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper reports mean and standard deviation across multiple runs with different random seeds and clearly states the sources of variability, ensuring the statistical reliability of the results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper specifies the type of GPUs used, training time per experiment, and overall compute requirements, providing sufficient details to assess reproducibility and resource demands.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research adheres to the NeurIPS Code of Ethics, with no identified ethical concerns related to data usage, human subjects, or societal impact.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper includes a Broader Impacts section stating that the work aims to advance machine learning and does not raise specific societal concerns requiring detailed discussion.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not introduce models or datasets with a high risk of misuse, so safeguards are not applicable.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and models used in the paper are properly cited, and their licenses and terms of use are respected and included where applicable.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release any new datasets, models, or code assets, so this question is not applicable.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects, so this question is not applicable.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects, so IRB approval is not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methods in this research do not involve LLMs in any important, original, or non-standard way, so this question is not applicable.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
A   Theoretical Risk Comparison ........................................................................................................................................................................24
B   Additional Descriptions ........................................................................................................................................................................26
  B.1  Details of Dataset and Task Settings ........................................................................................................................................................................26
  B.2  Details of Downstream Models ........................................................................................................................................................................27
  B.3  Details of Baselines ........................................................................................................................................................................27
  B.4  Details of Baseline Hyper-parameters ........................................................................................................................................................................29
  B.5  Comparison of Assumptions and Requirements ........................................................................................................................................................................29
C   Additional Results ........................................................................................................................................................................30
  C.1  Detailed Overall Performance Results ........................................................................................................................................................................30
  C.2  Accuracy Trends Across Sequential Tasks ........................................................................................................................................................................30
  C.3  Detailed Results Under Distribution Shifts ........................................................................................................................................................................30
  C.4  Inference Efficiency and Parameter Overhead ........................................................................................................................................................................34
  C.5  Forward Transfer Analysis ........................................................................................................................................................................34
  C.6  Additional Visualizations of Gate Activations and Relaxation Effect ........................................................................................................................................................................35
D   Discussions ........................................................................................................................................................................38
  D.1  Use of Unlabeled Adaptation Samples ........................................................................................................................................................................38
  D.1  Relation to Rehearsal-Free Continual Learning ........................................................................................................................................................................38
  D.3  Limitations ........................................................................................................................................................................38
  D.4  Broader Impacts ........................................................................................................................................................................38
Appendix A Theoretical Risk Comparison: Dynamic MoE vs. Static Averaging
Problem Setup and Definitions
Consider $T$ independent tasks, each associated with a data distribution $D_{t}$ for $t=1,\dots,T$. For each task $t$, a pre-trained expert model $f_{t}(x)$ outputs a probability distribution over classes, trained specifically on $D_{t}$. The overall data distribution $D$ is a mixture of these tasks, where an example $(x,y)$ is drawn from task $t$ with prior probability $P(t)$, and then $(x,y)\sim D_{t}$. The expected risk of a predictive model $h(x)$ is defined as:
$R(h)=\mathbb{E}_{(x,y)\sim D}\big[\ell(h(x),y)\big]=\sum_{t=1}^{T}P(t)\,\mathbb{E}_{(x,y)\sim D_{t}}\big[\ell(h(x),y)\big],$ (12)
where $\ell(h(x),y)$ is a loss function (e.g., cross-entropy or 0-1 loss).
We compare two methods to combine the experts into a final prediction $h(x)$:
  • Static Averaging: Defined as $h_{\text{static}}(x)=\sum_{i=1}^{T}\alpha_{i}f_{i}(x)$, where $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{T})$ is a fixed weight vector independent of $x$, typically with $\alpha_{i}\geq 0$ and $\sum_{i}\alpha_{i}=1$ for probability outputs.
  • Dynamic Mixture-of-Experts (MoE): Defined as $h_{\text{MoE}}(x)=f_{i^{*}(x)}(x)$, where $i^{*}(x)=\arg\max_{i}g_{i}(x)$ and $g(x)=(g_{1}(x),\dots,g_{T}(x))$ is a gating function that selects one expert per input (hard routing). The gating is subject to routing noise, modeled below.
Routing Noise Model
For each input $x$ drawn from $D_{t}$, the true task is $t$, and the ideal expert is $f_{t}$. The gating selects the correct expert $i^{*}(x)=t$ with probability $1-\varepsilon_{t}$, and an incorrect expert $i^{*}(x)\neq t$ with probability $\varepsilon_{t}=P(i^{*}(x)\neq t\mid x\sim D_{t})$, the task-specific routing error rate. On error, the gating selects a random expert from $\{1,\dots,T\}\setminus\{t\}$ uniformly. Define:
  • $R_{t}(i)=\mathbb{E}_{(x,y)\sim D_{t}}[\ell(f_{i}(x),y)]$, the risk of expert $i$ on task $t$.
  • $R_{\text{ideal}}=\sum_{t=1}^{T}P(t)R_{t}(t)$, the risk with perfect routing.
  • $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$, the average risk of incorrect experts on task $t$.
  • $\varepsilon=\sum_{t=1}^{T}P(t)\varepsilon_{t}$, the overall routing error rate.
Theorem A.1 (Dynamic MoE versus Static Averaging).
Let $\{(D_{t},f_{t})\}_{t=1}^{T}$ be $T$ independent tasks with priors $P(t)$ and per-task risks $R_{t}(i)$. For any static mixture $h_{\mathrm{static}}(x)=\sum_{i=1}^{T}\alpha_{i}\,f_{i}(x)$ and any hard-routed MoE $h_{\mathrm{MoE}}(x)=f_{i^{\star}(x)}(x)$ with task-specific routing errors $\varepsilon_{t}$: $R(h_{\mathrm{MoE}})\;=\;R_{\mathrm{ideal}}+\sum_{t=1}^{T}P(t)\,\varepsilon_{t}\bigl(R_{\text{wrong},t}-R_{t}(t)\bigr),$ (13) where $R_{\mathrm{ideal}}=\sum_{t}P(t)R_{t}(t)$ and $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$. Moreover, 1. (Perfect routing) If $\varepsilon_{t}=0$ for all $t$, then $\inf_{g}R(h_{\mathrm{MoE}})\;<\;\inf_{\boldsymbol{\alpha}}R(h_{\mathrm{static}})$ whenever at least two tasks disagree on their best expert. 2. (Noisy routing)  If $\sum_{t}P(t)\varepsilon_{t}\bigl(R_{\text{wrong},t}-R_{t}(t)\bigr)<R^{*}_{\mathrm{static}}-R_{\mathrm{ideal}},$ where $R^{*}_{\text{static}}=\inf_{\boldsymbol{\alpha}}R(h_{\text{static}})$, then the MoE still attains lower risk than any static mixture.
Proof.
The proof proceeds in three parts: (1) deriving the MoE risk with routing noise, (2) proving the optimal gating case, and (3) establishing the condition for MoE superiority under routing noise.
Step 1: MoE Risk with Routing Noise
The MoE prediction is $h_{\text{MoE}}(x)=f_{i^{*}(x)}(x)$, where $i^{*}(x)=\arg\max_{i}g_{i}(x)$. The expected risk is:
$R(h_{\text{MoE}})=\mathbb{E}_{(x,y)\sim D}[\ell(f_{i^{*}(x)}(x),y)]=\sum_{t=1}^{T}P(t)\,\mathbb{E}_{(x,y)\sim D_{t}}[\ell(f_{i^{*}(x)}(x),y)].$ (14)
For task $t$, condition on routing correctness:
  • Correct routing ($i^{*}(x)=t$): Probability $1-\varepsilon_{t}$, risk $R_{t}(t)$.
  • Incorrect routing ($i^{*}(x)\neq t$): Probability $\varepsilon_{t}$, selects a random expert from $\{1,\dots,T\}\setminus\{t\}$, with average risk $R_{\text{wrong},t}=\frac{1}{T-1}\sum_{i\neq t}R_{t}(i)$.
The expected risk on task $t$ is:
$\mathbb{E}_{(x,y)\sim D_{t}}[\ell(f_{i^{*}(x)}(x),y)]=(1-\varepsilon_{t})R_{t}(t)+\varepsilon_{t}R_{\text{wrong},t}.$ (15)
Thus, the total risk is:
$R(h_{\text{MoE}})=\sum_{t=1}^{T}P(t)\big[(1-\varepsilon_{t})R_{t}(t)+\varepsilon_{t}R_{\text{wrong},t}\big].$ (16)
Rewrite:
$R(h_{\text{MoE}})=\sum_{t=1}^{T}P(t)R_{t}(t)+\sum_{t=1}^{T}P(t)\varepsilon_{t}(R_{\text{wrong},t}-R_{t}(t))=R_{\text{ideal}}+\sum_{t=1}^{T}P(t)\varepsilon_{t}\delta_{t},$ (17)
where $\delta_{t}=R_{\text{wrong},t}-R_{t}(t)>0$ is the risk increase due to misrouting on task $t$.
Step 2: Optimal Gating (No Routing Noise)
Assume an oracle gating function with $\varepsilon_{t}=0$ for all $t$, so $i^{*}(x)=t$ for all $x\sim D_{t}$. Then:
$R(h_{\text{MoE}})=\sum_{t=1}^{T}P(t)R_{t}(t)=R_{\text{ideal}}.$ (18)
Define hypothesis classes:
$\mathcal{H}_{\text{static}}=\left\{x\mapsto\sum_{i=1}^{T}\alpha_{i}f_{i}(x)\mid\boldsymbol{\alpha}\in\mathbb{R}^{T}\right\},$ (19)
$\mathcal{H}_{\text{MoE}}=\left\{x\mapsto f_{i^{*}(x)}(x)\mid i^{*}(x)=\arg\max_{i}g_{i}(x),g:\mathcal{X}\to\mathbb{R}^{T}\right\}.$ (20)
Any static model $h_{\text{static}}(x)=\sum_{i=1}^{T}\alpha_{i}f_{i}(x)$ can be approximated by an MoE with $g(x)$ assigning constant weights, so $\mathcal{H}_{\text{static}}\subseteq\mathcal{H}_{\text{MoE}}$. Thus:
$R^{*}_{\text{MoE}}=\inf_{g(\cdot)}R(h_{\text{MoE}})\leq\inf_{\boldsymbol{\alpha}}R(h_{\text{static}})=R^{*}_{\text{static}}.$ (21)
Under task heterogeneity ($R_{t}(t)<R_{t}(s)$ and $R_{s}(s)<R_{s}(t)$ for some $t\neq s$), the ideal MoE routes each $x\sim D_{t}$ to $f_{t}$, achieving:
$R_{\text{ideal}}=\sum_{t=1}^{T}P(t)R_{t}(t).$ (22)
For static averaging:
$R(h_{\text{static}})=\sum_{t=1}^{T}P(t)\mathbb{E}_{(x,y)\sim D_{t}}\left[\ell\left(\sum_{i=1}^{T}\alpha_{i}f_{i}(x),y\right)\right].$ (23)
Since $\ell$ is convex (e.g., cross-entropy), Jensen’s inequality implies:
$\mathbb{E}_{D_{t}}\left[\ell\left(\sum_{i}\alpha_{i}f_{i}(x),y\right)\right]\geq R_{t}(t),$ (24)
with strict inequality unless $\alpha_{t}=1$ and $\alpha_{i}=0$ for $i\neq t$, which cannot hold for all tasks simultaneously under heterogeneity. Thus:
$R^{*}_{\text{static}}>R_{\text{ideal}}=R^{*}_{\text{MoE}}.$ (25)
Step 3: MoE Superiority with Routing Noise
Let $\gamma=R^{*}_{\text{static}}-R_{\text{ideal}}>0$ under task heterogeneity. The MoE outperforms the static model if:
$R(h_{\text{MoE}})<R^{*}_{\text{static}}.$ (26)
Substitute:
$R_{\text{ideal}}+\sum_{t=1}^{T}P(t)\varepsilon_{t}\delta_{t}<R_{\text{ideal}}+\gamma.$ (27)
Thus:
$\sum_{t=1}^{T}P(t)\varepsilon_{t}\delta_{t}<\gamma=R^{*}_{\text{static}}-R_{\text{ideal}}.$ (28)
Since $\delta_{t}=R_{\text{wrong},t}-R_{t}(t)$, the condition is:
$\sum_{t=1}^{T}P(t)\varepsilon_{t}(R_{\text{wrong},t}-R_{t}(t))<R^{*}_{\text{static}}-R_{\text{ideal}}.$ (29)
If this holds, the MoE’s risk, despite routing noise, remains below the best static risk. ∎
Conclusion: The MoE risk is $R_{\text{ideal}}+\sum_{t=1}^{T}P(t)\varepsilon_{t}(R_{\text{wrong},t}-R_{t}(t))$, and it outperforms static averaging when routing noise is sufficiently small relative to the static model’s suboptimality. The optimal gating case confirms $R^{*}_{\text{MoE}}\leq R^{*}_{\text{static}}$, with strict inequality under task heterogeneity.
Appendix B Additional Descriptions
B.1 Details of Dataset and Task Settings
Dataset Details
Following prior works [70], we evaluate continual model merging on twenty publicly available image classification datasets, including SUN397 [84], Stanford Cars [32], RESISC45 [4], EuroSAT [19], SVHN [46], GTSRB [64], MNIST [34], DTD [7], Flowers102 [47], PCAM [73], FER2013 [18], Oxford-IIIT Pet [50], STL-10 [9], CIFAR-100 and CIFAR-10 [33], Food-101 [3], Fashion-MNIST [83], EMNIST [10], KMNIST [8], and Rendered SST-2 [62].
Table 8: Extended downstream datasets used in our experiments.
Dataset #Classes #Train (k) #Test (k) Task
SUN397 287 19.9 19.9 Scene category
Stanford Cars 196 8.1 8.0 Car series
RESISC45 45 18.9 6.3 Remote–sensing scene
EuroSAT 10 21.6 2.7 Satellite land-use
SVHN 10 73.3 26.0 Digit recognition
GTSRB 43 39.2 12.6 Traffic sign
MNIST 10 60 10 Hand-written digit
DTD 47 3.8 1.9 Texture recognition
Flowers102 102 1.0 6.1 Flower species
PCAM 2 262 32.8 Tumour detection
FER2013 7 28.7 3.6 Facial emotion
Oxford IIIT Pet 37 3.7 3.7 Animal species
STL10 10 5 8 Object recognition
CIFAR-100 100 50 10 Natural object
CIFAR-10 10 50 10 Natural object
Food101 101 75.8 25.3 Food type
Fashion-MNIST 10 60 10 Fashion product
EMNIST (digits) 10 60 10 Hand-written digit
KMNIST 10 60 10 Kuzushiji character
Rendered SST-2 2 6.9 1.8 Rendered sentiment
Task Grouping
We group the 20 datasets into three progressive task sets and evaluate the merged models using average accuracy (ACC) and backward transfer (BWT) metrics. For each task group, we perform 10 experiments using different task sequences (listed in Tab.  9), and report both the mean and standard deviation of the results to ensure robustness and consistency.
  • 8-task group: (1) SUN397, (2) Stanford Cars, (3) RESISC45, (4) EuroSAT, (5) SVHN, (6) GTSRB, (7) MNIST, (8) DTD.
  • 14-task group: (1) SUN397, (2) Stanford Cars, (3) RESISC45, (4) EuroSAT, (5) SVHN, (6) GTSRB, (7) MNIST, (8) DTD, (9) Flowers102, (10) PCAM, (11) FER2013, (12) OxfordIIITPet, (13) STL10, (14) CIFAR100.
  • 20-task group: (1) SUN397, (2) Stanford Cars, (3) RESISC45, (4) EuroSAT, (5) SVHN, (6) GTSRB, (7) MNIST, (8) DTD, (9) Flowers102, (10) PCAM, (11) FER2013, (12) OxfordIIITPet, (13) STL10, (14) CIFAR100, (15) CIFAR10, (16) Food101, (17) FashionMNIST, (18) EMNIST, (19) KMNIST, (20) RenderedSST2.
Table 9: Dataset orderings used for experiments in each task group.
Group Order Dataset Order (by ID)
[…]
B.2 Details of Downstream Models
In this section, we present the evaluation setup for pre-trained and fine-tuned models. As shown in Tab. 10, we evaluate the zero-shot accuracy of the original CLIP-ViT models and the performance of fine-tuned models on the test sets of various downstream tasks. The fine-tuned checkpoints are obtained directly from Hugging Face (https://huggingface.co/tanganke), where each model has been fine-tuned on task-specific training data using a standard protocol. The visual encoder is updated during fine-tuning, while the classification head is fixed and initialized from the pre-trained text encoder. The fine-tuning setup follows a standard configuration: cross-entropy loss, Adam optimizer, cosine annealing learning rate schedule with a peak learning rate of 1e-5, batch size 128, and 4000 training steps.
Table 10: Test set accuracy of the pre-trained model and individual fine-tuned models on different downstream tasks.
Model
SUN397
Cars
RESISC45
EuroSAT
SVHN
GTSRB
MNIST
DTD
Flowers102
PCAM
CLIP-ViT-B/32
Pre-trained 63.2 59.6 60.3 45.0 31.6 32.5 48.3 44.2 66.4 60.6
Fine-tuned 74.9 78.5 95.1 99.1 97.3 98.9 99.6 79.7 88.6 88.0
CLIP-ViT-B/16
Pre-trained 65.5 64.7 66.4 54.1 52.0 43.5 51.7 45.0 71.3 54.0
Fine-tuned 78.9 85.9 96.6 99.0 97.6 99.0 99.7 82.3 94.9 90.6
CLIP-ViT-L/14
Pre-trained 68.2 77.9 71.3 61.2 58.4 50.5 76.3 55.5 79.2 51.2
Fine-tuned 82.8 92.8 97.4 99.1 97.9 99.2 99.8 85.5 97.7 91.1
Model
FER2013
OxfordIIITPet
STL10
CIFAR100
CIFAR10
Food101
FashionMNIST
EMNIST
KMNIST
RenderedSST2
CLIP-ViT-B/32
Pre-trained 41.3 83.3 97.1 63.7 89.8 82.4 63.0 12.0 10.0 58.6
Fine-tuned 71.6 92.5 97.5 88.4 97.6 88.4 94.7 95.6 98.2 71.3
CLIP-ViT-B/16
Pre-trained 46.4 88.4 98.3 66.3 90.8 87.0 67.3 12.4 11.2 60.6
Fine-tuned 72.8 94.5 98.2 88.8 98.3 91.9 94.5 95.3 98.1 75.7
CLIP-ViT-L/14
Pre-trained 50.0 93.2 99.4 75.1 95.6 91.2 67.0 12.3 9.7 68.9
Fine-tuned 75.9 95.7 99.2 93.0 99.1 94.8 95.3 95.4 98.3 80.5
B.3 Details of Baselines
Our experiments involve the following comparison methods and our method:
  • Stochastic Weight Averaging (SWA). A simple model averaging technique to stabilize optimization and improve generalization [26]. At each step $t$, the model parameters are averaged across previous checkpoint: $\theta^{\text{SWA}}_{t}=\frac{1}{t}\left[\theta^{\text{SWA}}_{t-1}(t-1)+\theta^{\text{SWA}}_{t}\right]$. This approach can be interpreted as a form of uniform model ensembling. While conceptually straightforward, SWA treats all checkpoints equally and does not account for inter-task conflicts.
  • Continual Task Arithmetic (C. TA). A training-free merging strategy that linearly combines task-specific fine-tuned models with a shared pre-trained model [24]. It computes the merged parameters as $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda(\theta_{t}-\theta_{0})$, where $\lambda$ is a scaling factor. TA is computationally efficient and easy to apply, but sensitive to $\lambda$ and prone to destructive interference when merging dissimilar tasks.
  • Continual Ties-Merging (C. Ties). An extension of Task Arithmetic that reduces parameter-level redundancy and sign conflicts during model merging [86]. For task $t$, the difference vector $\Delta\theta_{t}=\theta_{t}-\theta_{0}$ is trimmed and sign-normalized to obtain $\Delta\theta^{\text{Ties}}_{t}=\mathrm{Ties}{\left(\Delta\theta^{\text{Ties}}_{t-1},\Delta\theta_{t}\right)}$, and the merged model is given by $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda\Delta\theta^{\text{Ties}}_{t}$.
  • Orthogonal Projection-based Continual Merging (OPCM). A projection-based scheme to mitigate task interference by enforcing orthogonality between parameter updates [70]. Specifically, each $\Delta\theta_{t}$ is projected onto the orthogonal complement of the subspace spanned by previous updates: $\theta^{\text{merged}}_{t}=\theta_{0}+\frac{1}{\lambda_{t}}\left[\lambda_{t-1}\Delta\theta^{\text{merged}}_{t-1}+\mathcal{P}^{(t-1)}(\Delta\theta_{t})\right]$, where $\mathcal{P}^{(t-1)}$ denotes the orthogonal projection.
  • Maximum Magnitude Selection (MagMax). An extension of Task Arithmetic that, for each parameter dimension, selects the update with the larger absolute value: $\Delta\theta^{\text{MagMax}}_{t}=\mathrm{MagMax}{\left(\Delta\theta^{\text{MagMax}}_{t-1},\Delta\theta_{t}\right)}$, and the merged model is given by $\theta^{\text{merged}}_{t}=\theta^{\text{merged}}_{t-1}+\lambda\Delta\theta^{\text{MagMax}}_{t}$.
B.4 Details of Baseline Hyper-parameters
Tab. 11 summarizes the hyper-parameters for all baseline methods under different task configurations (8, 14, 20 tasks). Top-k denotes the pruning ratio, TALL the TALL mask threshold, and Cons. the consensus mask threshold. The column LR is the learning rate, while Steps indicates the number of adaptation steps. $r$ represents the LoRA rank, and the last column jointly reports the null dimension ($k$), EMA decay ($\beta$), and relaxation coefficient ($\gamma$).
Table 11: hyper-parameter settings for all baselines.
Method Tasks $\lambda$ Top-k TALL Cons. LR Steps $r$ $k/\beta/\gamma$
Task Arithmetic 8 0.3 - - - - - - -
14/20 0.1 - - - - - - -
Ties-Merging 8 0.3 20 - - - - - -
14/20 0.1 20 - - - - - -
Consensus TA 8/14/20 0.1 - 0.2 2 - - - -
LW. AdaMerging 8/14/20 0.3 - - - 1e-4 50 - -
WEMOE 8/14/20 0.3 - - - 1e-4 50 - -
LoRA-WEMOE 8/14/20 0.3 - - - 1e-4 50 64 -
MINGLE-Static 8/14/20 0.3 - - - - - - -
Mingle (Ours) 8/14/20 - - - - 1e-4 50 64 3/0.99/1.0
B.5 Comparison of Assumptions and Requirements
Tab. 12 summarizes the assumptions and resource requirements of all baseline methods. We report whether each method requires storing intermediate activations, introduces additional parameters (for storage or inference), and incurs extra test-time computation. Our method only maintains a fixed-size covariance matrix instead of full activations, leading to constant memory regardless of test set size. Although LoRA experts are stored, the router merges them into a single model per input, so the effective inference cost matches that of a standard individual model.
Table 12: Comparison of baseline assumptions and requirements.
Method Save Activations Extra Parameters Extra Parameters Test-time Compute
(Storage) (Inference)
Task Arithmetic No No No No
Ties-Merging No No No No
MAGMAX-Ind No No No No
OPCM No No No No
Consensus TA No Yes No No
LW. AdaMerging No No No Yes
WEMOE No Yes No Yes
MINGLE-Static No No No No
Mingle (Ours) No1 Yes2 No2 Yes
1 Only a fixed-size covariance matrix is maintained, resulting in constant memory regardless of test set size.
2 LoRA experts are stored, but the router merges them into a single model per input, making the effective inference size equivalent to a standard individual model.
Appendix C Additional Results
In this section, we provide additional experimental results to support the findings reported in the main paper. Specifically, we include: (1) detailed overall performance results (C.1); (2) accuracy trends across sequential tasks (C.2); (3) detailed results under distribution shifts (C.3); and (4) extended visualizations of gate activations and hyper-parameter effects (C.6).
C.1 Detailed Overall Performance Results
Tab. 13 expands on the average results in Tab. 1 by reporting per-task average accuracy after continually merging 20 tasks. We compare six methods, SWA, Task Arithmetic, Ties-Merging, MagMax-IND, OPCM, and our proposed Mingle across three CLIP-ViT backbones (B/32, B/16, L/14). Mingle achieves the highest accuracy on most tasks. These fine-grained results reinforce the main paper’s findings, highlighting Mingle’s ability to improve performance on continual model merging.
Table 13: Test set accuracy comparisons on different downstream tasks.
Model
SUN397
Cars
RESISC45
EuroSAT
SVHN
GTSRB
MNIST
DTD
Flowers102
PCAM
ViT-B/32
C. Fine-Tuned 53.9 38.2 64.7 98.7 45.4 34.4 86.7 58.4 57.5 67.7
Average (SWA) 64.2 59.6 64.8 60.9 47.3 43.1 71.8 46.4 66.5 63.9
C.TA 62.0 53.7 60.9 58.1 48.5 48.9 79.4 46.1 61.1 73.4
C.TIES 62.5 49.1 55.8 50.9 54.6 49.3 82.0 46.7 58.5 69.9
MagMax-Ind 63.6 53.1 59.7 49.1 53.8 53.1 79.8 43.2 56.9 75.1
Consensus TA 37.0 25.2 35.2 36.7 37.3 44.1 80.6 30.3 33.5 59.2
C. LW AdaMerging 63.1 60.0 63.5 60.1 35.6 32.1 51.8 45.4 66.6 60.2
C. LoRA-WEMOE 51.4 45.8 63.3 43.5 42.9 34.6 58.9 46.5 47.5 60.1
OPCM 64.4 51.1 66.0 71.7 66.1 56.0 90.2 40.4 64.9 80.2
Mingle (Ours) 67.8 58.3 83.5 90.0 82.9 91.8 98.0 65.3 74.0 66.9
Mingle (Ours) 68.8 64.2 83.8 91.1 82.4 89.0 96.9 62.8 76.7 72.8
ViT-B/16
C. Fine-Tuned 62.7 58.0 67.6 99.1 46.0 29.2 93.9 61.9 64.1 75.2
Average (SWA) 67.1 64.6 69.3 63.4 62.4 52.7 80.7 46.6 71.8 63.1
C.TA 65.8 57.5 63.8 59.5 64.7 54.0 88.0 45.3 67.5 67.1
C.TIES 64.2 52.9 60.9 53.0 62.8 48.8 88.4 45.0 61.3 68.5
MagMax-Ind 65.8 51.8 57.8 42.6 54.4 43.7 83.0 42.8 60.4 69.8
Consensus TA 42.6 24.8 30.4 34.4 47.6 42.2 79.9 30.6 36.2 74.3
C. LW AdaMerging 65.5 65.7 69.8 59.4 50.1 44.2 61.1 47.1 71.8 57.9
C. LoRA-WEMOE 62.7 60.2 69.4 37.7 52.1 39.9 63.1 45.3 64.3 51.7
OPCM 67.9 55.9 73.7 77.5 74.4 63.2 94.1 49.2 72.3 79.6
Mingle (Ours) 71.5 64.9 85.3 90.0 87.5 90.1 97.1 62.7 82.6 80.6
Mingle (Ours) 72.0 72.1 87.9 93.3 87.1 89.2 97.4 62.5 86.8 76.4
ViT-L/14
C. Fine-Tuned 69.5 73.6 78.3 99.2 59.3 49.3 98.6 69.7 83.2 78.3
Average (SWA) 70.7 77.7 76.4 75.3 69.5 62.1 93.7 57.7 80.0 73.6
C.TA 70.4 74.1 73.9 66.3 69.9 65.6 95.1 56.6 78.6 70.4
C.TIES 69.7 70.3 65.3 47.9 76.1 63.6 94.7 54.4 77.9 72.3
MagMax-Ind 73.1 73.7 75.6 64.6 73.7 68.8 94.6 56.1 78.0 71.7
Consensus TA 50.7 39.1 31.7 36.4 39.4 44.9 88.5 33.8 45.7 62.5
C. LW AdaMerging 68.8 78.6 75.9 65.7 58.3 51.6 79.9 57.4 80.6 52.4
C. LoRA-WEMOE 62.1 68.1 68.7 53.2 47.5 49.4 69.8 49.1 66.2 54.2
OPCM 73.1 78.3 82.4 80.2 80.8 80.4 97.4 61.6 84.8 76.3
Mingle (Ours) 75.9 83.4 87.8 88.7 91.1 94.5 98.4 70.8 94.8 75.3
Mingle (Ours) 74.5 85.9 90.5 92.5 90.1 92.7 98.1 69.2 95.7 74.0
Model
FER2013
OxfordIIITPet
STL10
CIFAR100
CIFAR10
Food101
FashionMNIST
EMNIST
KMNIST
RenderedSST2
ViT-B/32
C. Fine-Tuned 58.3 68.5 86.7 40.2 70.5 50.0 90.7 72.4 54.5 54.5
Average (SWA) 50.2 84.1 97.0 69.8 92.7 80.4 71.3 15.0 11.5 61.8
C.TA 51.4 82.3 94.9 64.6 91.4 71.9 73.9 17.8 12.2 59.9
C.TIES 49.5 81.3 95.2 63.7 91.2 70.2 73.7 17.8 16.9 59.8
MagMax-Ind 56.5 79.9 94.6 68.7 91.9 73.8 74.3 18.3 15.4 63.9
Consensus TA 41.7 58.8 81.8 41.5 78.1 29.8 72.6 17.4 18.5 54.1
C. LW AdaMerging 43.2 83.7 96.8 67.0 89.9 81.6 63.7 16.8 10.7 59.1
C. LoRA-WEMOE 44.6 72.5 86.1 40.1 63.8 63.8 48.1 10.3 12.8 55.7
OPCM 58.5 82.9 95.9 67.6 92.8 74.0 76.3 22.4 18.3 64.6
Mingle (Ours) 65.0 85.5 97.0 72.6 94.1 81.5 85.4 50.4 65.2 67.1
Mingle (Ours) 65.3 88.5 97.7 73.9 94.7 83.7 86.4 39.3 56.1 68.7
ViT-B/16
C. Fine-Tuned 60.5 84.5 90.5 38.8 73.6 61.9 89.7 83.3 51.5 72.8
Average (SWA) 50.9 89.6 98.0 72.9 94.2 85.9 73.3 15.6 12.4 62.5
C.TA 50.7 89.3 97.0 68.0 93.1 80.3 75.7 18.1 16.7 61.8
C.TIES 50.4 87.9 96.3 63.1 91.7 78.0 75.0 23.4 24.9 61.5
MagMax-Ind 57.7 88.8 97.5 71.5 94.4 81.3 77.2 24.5 25.0 59.4
Consensus TA 45.6 76.8 87.7 44.4 82.2 38.4 72.7 18.8 30.0 58.6
C. LW AdaMerging 46.8 88.9 98.1 69.2 91.4 86.6 67.2 17.2 11.0 59.2
C. LoRA-WEMOE 45.6 91.2 92.3 41.3 64.3 78.1 48.0 23.5 16.6 52.7
OPCM 59.5 91.8 97.7 73.2 94.7 83.1 81.3 26.5 23.4 66.8
Mingle (Ours) 67.6 92.7 97.4 74.0 95.3 87.7 87.4 73.5 79.9 74.0
Mingle (Ours) 67.9 93.5 98.4 77.7 96.4 89.7 87.8 56.6 64.5 75.3
ViT-L/14
C. Fine-Tuned 68.0 92.1 94.5 60.5 85.7 74.8 93.1 89.0 59.2 78.8
Average (SWA) 52.7 94.2 99.2 81.7 97.0 90.7 77.4 16.1 10.4 66.1
C.TA 55.7 94.2 98.6 79.1 96.6 87.6 80.8 17.6 10.6 63.6
C.TIES 57.6 93.5 97.8 74.0 95.6 84.7 79.7 20.2 12.6 58.4
MagMax-Ind 52.9 93.9 98.7 82.1 97.3 89.5 81.6 19.2 11.1 68.4
Consensus TA 50.3 82.2 89.7 47.5 86.2 43.5 75.3 14.5 10.4 53.4
C. LW AdaMerging 49.2 93.5 99.3 77.2 95.8 91.1 68.2 18.6 9.8 66.6
C. LoRA-WEMOE 46.3 84.5 87.6 52.1 70.5 73.3 50.0 18.7 10.9 56.5
OPCM 61.8 95.4 99.2 83.0 97.8 90.9 86.0 26.4 14.7 71.0
Mingle (Ours) 67.7 96.0 98.7 81.4 97.1 90.6 90.6 60.7 88.6 79.8
Mingle (Ours) 67.9 96.0 99.4 84.7 97.8 92.4 88.8 53.0 57.1 75.5
C.2 Accuracy Trends Across Sequential Tasks
Fig. 5 provides a detailed view of accuracy throughout the continual merging process across different settings, showing both the performance on the current task and on previously encountered ones. The progressive accuracy drop across columns illustrates the degree of forgetting over time. Notably, Mingle consistently alleviates this degradation, demonstrating markedly reduced forgetting across the full task sequence. Fig. 6 further compares the average accuracy curves of Mingle and baseline methods on previously seen tasks after each new model is merged, using the CLIP ViT-B/16 backbone. Results are averaged over 10 random task orderings. Mingle consistently achieves the highest performance throughout the merging process, with its accuracy curve clearly dominating those of competing methods. Moreover, the narrower standard deviation bands indicate that Mingle is more robust to the task orders.
C.3 Detailed Results Under Distribution Shifts
Tab. 14 expands on Tab. 4 by reporting per-dataset accuracy under both clean test conditions and seven common corruption types: motion blur, impulse noise, Gaussian noise, pixelation, spatter, contrast shift, and JPEG compression. We evaluate six merging methods, across four downstream tasks: Cars, EuroSAT, RESISC45, and GTSRB. This detailed breakdown complements the average results in the main paper, providing a more comprehensive assessment of robustness under test-time distribution shifts.
(a) ViT-B/32 on 8 Tasks
(b) ViT-B/16 on 8 Tasks
(c) ViT-L/14 on 8 Tasks
(d) ViT-B/16 on 14 Tasks
(e) ViT-L/14 on 14 Tasks
(f) ViT-L/14 on 20 Tasks
Figure 5: Accuracy matrices of Mingle (ViT-B/32, ViT-B/16, and ViT-L/14) under different task settings.
(a) 8 tasks
(b) 14 tasks
(c) 20 tasks
Figure 6: Sequential test accuracy curves of Mingle and baselines (C.TA, C.Ties, MagMax, OPCM) under different task settings. Shaded regions indicate standard deviation across 10 task orders.
Table 14: Robustness results when merging ViT-B/32 models on four tasks.
Clean Test Set Corruption: Motion Blur
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 65.3 49.7 65.4 43.8 56.0 64.2 25.6 62.5 37.5 47.5
C. WEMOE 0.5 8.1 2.6 2.5 3.4 0.5 8.0 1.8 2.3 3.1
C. LoRA-WEMOE 66.1 84.3 81.0 83.6 78.7 64.8 57.9 82.0 79.2 71.0
C. Task Arithmetic 64.6 90.4 80.2 74.8 77.5 62.3 59.4 78.5 63.3 65.9
MagMax-Ind 63.1 89.2 81.7 82.5 79.1 61.4 62.1 80.0 72.6 69.0
OPCM 65.7 92.3 85.7 90.5 83.6 62.8 62.5 83.7 82.2 72.8
Mingle (Ours) 74.4 96.5 91.5 97.3 89.9 73.2 70.5 91.9 95.8 82.9
Corruption: Impulse Noise Corruption: Gaussian Noise
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 60.5 30.1 56.3 25.5 43.1 62.3 25.6 59.7 25.6 43.3
C. WEMOE 0.5 11.2 2.3 3.2 4.3 0.5 8.1 2.4 2.8 3.4
C. LoRA-WEMOE 62.2 23.4 69.9 64.6 55.0 64.9 31.7 77.8 63.4 59.4
C. Task Arithmetic 59.9 57.7 72.9 45.0 58.9 61.8 51.4 75.1 50.1 59.6
MagMax-Ind 59.2 56.3 74.3 52.5 60.6 60.6 51.7 77.0 56.5 61.5
OPCM 61.1 57.1 78.5 62.0 64.7 63.0 52.4 80.7 64.9 65.2
Mingle (Ours) 69.6 28.0 86.1 86.1 67.5 72.0 38.5 89.4 82.9 70.7
Corruption: Pixelate Corruption: Spatter
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 3.4 16.5 13.5 39.2 18.1 61.3 34.1 58.2 32.8 46.6
C. WEMOE 0.5 6.3 2.5 2.5 3.0 0.5 10.1 2.7 2.6 4.0
C. LoRA-WEMOE 0.8 26.0 5.8 67.0 24.9 62.4 35.4 71.2 73.0 60.5
C. Task Arithmetic 2.5 31.7 19.1 65.6 29.7 61.2 63.1 72.7 57.0 63.5
MagMax-Ind 2.6 36.1 19.3 74.0 33.0 60.0 64.9 74.8 66.1 66.4
OPCM 2.1 34.3 19.5 84.9 35.2 61.5 64.7 78.8 76.8 70.5
Mingle (Ours) 2.3 35.6 18.5 95.1 37.9 70.1 57.8 86.2 93.9 77.0
Corruption: Contrast Corruption: JPEG Compression
Method Cars EuroSAT RESISC45 GTSRB Avg ACC Cars EuroSAT RESISC45 GTSRB Avg ACC
C. LW AdaMerging 61.8 26.0 63.1 44.8 48.9 65.1 29.6 65.4 36.4 49.1
C. WEMOE 0.5 7.5 2.3 3.0 3.3 0.5 10.5 2.4 2.7 4.0
C. LoRA-WEMOE 64.3 46.5 77.7 85.6 68.5 65.5 59.1 80.4 74.0 69.7
C. Task Arithmetic 62.5 55.2 75.3 70.8 66.0 64.1 66.2 80.0 61.0 67.8
MagMax-Ind 61.3 58.0 76.9 78.2 68.6 62.5 67.7 81.1 68.5 69.9
OPCM 63.8 57.5 81.3 87.4 72.5 65.0 68.0 85.4 79.3 74.4
Mingle (Ours) 72.4 60.1 90.4 97.3 80.1 73.7 73.5 92.0 92.4 82.9
C.4 Inference Efficiency and Parameter Overhead
Tab. 15 compares the inference efficiency and parameter overhead of all baselines on the CLIP ViT-B/32 model after merging eight tasks. We report the total number of parameters, additional storage and inference parameters, throughput (images per second), and accuracy. The results show that most static merging methods (e.g., Task Arithmetic, Ties-Merging, MAGMAX-Ind, OPCM) incur no extra storage or inference overhead, but typically achieve limited accuracy. Consensus TA and WEMOE introduce significant storage overhead, while WEMOE also scales up inference parameters considerably. By contrast, Mingle achieves a favorable trade-off: although it introduces additional parameters for LoRA experts and the router, the effective inference overhead remains small, and throughput is only marginally reduced compared to static baselines. This efficiency advantage comes while delivering substantially higher accuracy.
Table 15: Comparison of inference efficiency and parameter overhead on CLIP ViT-B/32 model after eight tasks merging.
Method Total Params (M) Extra Storage (M) Extra Inference (M) Throughput (img/s) ACC (%)
Task Arithmetic 87.5 0.0 0.0 $\sim$910 67.5
Ties-Merging 87.5 0.0 0.0 $\sim$910 49.0
MAGMAX-Ind 87.5 0.0 0.0 $\sim$910 70.7
OPCM 87.5 0.0 0.0 $\sim$910 75.5
Consensus TA 87.5 87.5 0.0 $\sim$910 69.0
LW. AdaMerging 87.5 0.0 0.0 $\sim$910 52.9
WEMOE 540.9 453.4 0.07 $\sim$858 4.9
WEMOE-LoRA 103.7 16.2 0.07 $\sim$848 66.6
MINGLE 173.1 85.6 0.6 $\sim$841 85.8
MINGLE* 113.7 26.2 0.3 $\sim$862 85.0
Table 16: Forward transfer (FWT) results on 8-task continual merging with CLIP ViT-B/16.
Method ACC (%) BWT (%) FWT (%)
Task Arithmetic 77.1 $\pm$ 0.0 -4.2 $\pm$ 1.0 -13.4 $\pm$ 0.0
Ties-Merging 66.8 $\pm$ 3.7 -5.5 $\pm$ 0.4 -30.7 $\pm$ 9.9
OPCM 81.8 $\pm$ 0.3 -4.8 $\pm$ 0.7 -9.0 $\pm$ 0.4
Mingle (Ours) 88.3 $\pm$ 0.6 -0.4 $\pm$ 0.1 -3.8 $\pm$ 0.8
C.5 Forward Transfer Analysis
Forward transfer (FWT) is an important metric in continual learning, as it quantifies how effectively prior knowledge facilitates the learning of future tasks. We adopt the standard definition:
$\text{FWT}=\frac{1}{T-1}\sum_{t=2}^{T}\left[a_{t}(\theta_{t}^{\text{merged}})-\bar{a}_{t}\right],$ (30)
where $a_{t}(\theta_{t}^{\text{merged}})$ denotes the test accuracy on task $t$ using the merged model after task $t$, and $\bar{a}_{t}$ is the accuracy of the individually fine-tuned model for task $t$. Positive FWT indicates beneficial transfer, while negative values suggest interference.
Tab. 16 reports the results for the 8-task continual merging setup on CLIP ViT-B/16. The results demonstrate that Mingle achieves nearly zero forgetting (BWT $\approx 0$) while obtaining the highest forward transfer among all baselines, showing that our adaptive gating and merging strategy not only preserves past knowledge but also enhances feature utility for future tasks.
C.6 Additional Visualizations of Gate Activations and the Relaxation Effect
We provide an extended ablation study on gate hyper-parameters, including visualizations of gate activations under 14-task (Fig. 7) and 20-task (Fig. 8) configurations, complementing the 8-task results presented in the main paper. The visualizations demonstrate that the null-space constraint remains effective as the number of tasks increases, consistently suppressing gate responses to inputs from previously seen tasks and thereby mitigating forgetting.
Figure 7: Visualization of gate activations across 14 tasks under varying $\gamma$ values. Each subplot corresponds to a gate, with curves and shaded regions denoting the mean and standard deviation of activations across layers. Gray bars mark the training dataset for each gate. Smaller $\gamma$ values result in stronger suppression of activations on previously learned tasks.
Figure 8: Visualization of gate activations across 20 tasks under varying $\gamma$ values.
Appendix D Discussions
D.1 Use of Unlabeled Adaptation Samples
In our experiments, we simulate a realistic deployment setting by randomly sampling 5 unlabeled examples per class from the test split, which serve as adaptation samples for model merging. Such small unlabeled buffers are practical in real-world applications and can be obtained from various sources, including (i) incoming test-time data such as recent user queries or model inputs, (ii) held-out validation inputs or small training subsets (if available), (iii) user-provided samples (e.g., few-shot examples) that do not raise privacy concerns, (iv) synthetically generated data, or (v) manually curated public data. Importantly, our method does not depend on precise sample selection, and the buffer size remains fixed and small, ensuring feasibility and robustness in deployment scenarios.
D.2 Relation to Rehearsal-Free Continual Learning
Test-time continual model merging (TTCMM) is closely related to the paradigm of rehearsal-free continual learning (RFCL), as both approaches share two fundamental constraints: (i) they do not retain past training data, and (ii) they avoid storing previous task models. The key distinction lies in the information available at each stage. RFCL assumes access to the training data of the current task and incrementally fine-tunes a single model over time. In contrast, TTCMM assumes access to independently fine-tuned models for each new task and focuses on merging these expert models rather than training them from scratch. Additionally, TTCMM relies on a small unlabeled buffer at test time (e.g., 5 samples per class) to guide the merging process.
From a privacy perspective, TTCMM provides stronger guarantees. Since it does not require access to full training sets, it only depends on a small set of unlabeled samples, which can be user-provided without risk, synthetically generated, or curated from public data. By comparison, RFCL requires access to large-scale labeled datasets for every task, raising more significant concerns regarding privacy, storage, and legal constraints (e.g., medical images, personal data, or copyrighted corpora). The reliance on a tiny unlabeled buffer makes TTCMM more practical in scenarios where data privacy is a primary consideration.
D.3 Limitations
As with many model merging methods, our approach assumes that all independently fine-tuned models originate from a shared pretrained initialization. The extent to which this assumption influences merging performance remains unclear and warrants further investigation. In addition, our current experiments focus on merging models with identical backbone architectures (e.g., CLIP ViT-B/16). Although our use of LoRA-based expert offers some structural uniformity, which could potentially accommodate heterogeneous backbones, we have not yet explored this setting. Extending our framework to support diverse initialization points or architectural variants remains an open direction for future work.
D.4 Broader Impacts
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.